### v1

# Characterising 1-Rectifiable metric spaces via Connected Tangent Spaces

Abstract.
We prove that in a complete metric space $X$, $1$-rectifiability of a set $E\subset X$ with $\mathcal{H}^{1}(E)<\infty$ and positive lower density $\mathcal{H}^{1}$-a.e. is implied by the property that all tangent spaces are connected metric spaces.
David Bate is supported by the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 948021). Phoebe Valentine is supported by the Warwick Mathematics Institute Centre for Doctoral Training, and gratefully acknowledges funding from the University of Warwick and the UK Engineering and Physical Sciences Research Council (Grant number: EP/W524645/1)
1. Introduction
A central concept in geometric measure theory is that of rectifiability. For a metric space $(X,d)$, a set $E\subset X$ is $n$-rectifiable if there exists countably many Lipschitz maps $f_{i}\colon A_{i}\subset\mathbb{R}^{n}\rightarrow X$ such that
$\mathcal{H}^{n}(E\backslash\bigcup_{i\in\mathbb{N}}f_{i}(A_{i}))=0.$
Here and throughout, $\mathcal{H}^{n}$ denotes the $n$-dimensional Hausdorff (outer) measure. A set is purely $n$-unrectifiable if its intersection with any $n$-rectifiable set has $\mathcal{H}^{n}$ measure zero.
This paper concerns characterizations of rectifiable metric spaces in terms of tangent spaces. In classical geometric measure theory, that is, for subsets of Euclidean space, Preiss [11] established a theory of tangent measures. Further, the rectifiability criterion of Marstrand [8] and Mattila [9] shows that a set $E\subset\mathbb{R}^{m}$ (with $\mathcal{H}^{n}(E)<\infty$ and positive lower density almost everywhere) is $n$-rectifiable if, for $\mathcal{H}^{n}$-a.e. $x\in E$, all tangent measures of $\mathcal{H}^{n}|_{E}$ at $x$ are supported on $n$-dimensional subspace of $\mathbb{R}^{m}$.
In the setting of metric spaces, the work of Kirchheim [5] shows that rectifiable sets possess a strong tangent structure. In particular, this work implies that tangent spaces (defined as a suitable Gromov–Hausdorff type adaptation of tangent measures, see Section 2) of a $n$-rectifiable set are Lebesgue measure supported on a normed $\mathbb{R}^{n}$. It remained an open question for sometime whether a result analogous to the Marstrand–Mattila theorem holds in metric spaces. Recently, this was shown to be the case. Indeed, Bate [2] shows that a metric space $E$ (with $\mathcal{H}^{n}$) is $n$-rectifiable whenever, for $\mathcal{H}^{n}$-a.e. $x\in E$, all tangent spaces of $\mathcal{H}^{n}|_{E}$ at $x$ are supported on biLipschitz images of $\mathbb{R}^{n}$.
In this article we show that, for 1-dimensional sets, this sufficient condition for rectifiability can be weakened further to tangents that are supported on connected metric spaces.
Let $\mathcal{C}_{*}$ be the space of pointed metric measure spaces $(X,d,\mu,x)$ for which the support of $\mu$ is connected.
Theorem 1.1.
Let $(X,d)$ be a complete metric space and let $E\subset X$ be $\mathcal{H}^{1}$-measurable with $\mathcal{H}^{1}(E)<\infty$. The following are equivalent:
  1. (1)
    $E$ is $1$-rectifiable.
  2. (2)
    For $\mathcal{H}^{1}$-a.e. $x\in E$, $\Theta_{*}^{1}(E,x)>0$ and
    $\textnormal{Tan}(X,d,\mathcal{H}^{1}\llcorner E,x)=\{(\mathbb{R},|\cdot|,\mathcal{H}^{1}/2,0)\}.$
  3. (3)
    For $\mathcal{H}^{1}$-a.e. $x\in E$, $\Theta_{*}^{1}(E,x)>0$ and
    $\textnormal{Tan}(X,d,\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{C}^{*}.$
The fact that 1) implies 2) is given in [2, Theorem 6.6] by applying the results of Kirchheim [5]. Since 2) implies 3) is immediate, our attention is focused on the implication 3) implies 1). We emphasise that, even in the case that all tangent spaces are supported on $\mathbb{R}$, our proof follows a completely different approach that of [2].
Indeed, the geometric idea of our proof begins with a result o Besicovitch [3, Lemma 9] that covers a purely 1-unrectifiable subset of $\mathbb{R}^{2}$ by a large number of circle pairs. A circle pair of a set $F$ is simply $B(x,|x-y|)\cup B(y,|x-y|)$, where $x$ and $y$ are points in $F$. Besicovitch demonstrates that for a $1$-unrectifiable set $F\subset\mathbb{R}^{2}$ and any $r>0$, one may find disjoint circle pairs, whose radii are at most $r$ and sum to at least $\mathcal{H}^{1}(F)/9$, such that the common regions are disjoint from $F$:
$F\cap B(x,|x-y|)\cap B(y,|x-y|)=\emptyset.$
This construction can be used to show that $1$-unrectifiable sets in $\mathbb{R}^{2}$ (with finite Hausdorff measure and positive lower density) cannot have tangent measures that are supported on lines: any such tangent is forced to be close to both $x$ and $y$ and hence must cross the empty region, thereby contradicting the idea that a tangent must stay close to the set it is approximating.
There are two issues in extending this idea to prove Theorem 1.1. Firstly, his proof only works in $\mathbb{R}^{2}$, in particular it uses the fact that boundaries of balls are continua. Secondly, when considering tangents supported on connected sets (even in $\mathbb{R}^{2}$), Besicovitch’s circle pairs are no longer sufficient: a connected tangent can lie close to both $x$ and $y$ and wander outside $B(x,|x-y|)\cup B(y,|x-y|)$ and thereby avoid the common region.
To solve this, we strengthen Besicovitch’s notion of circle pairs to what we term Besicovitch partitions, see Definition 4.5. Roughly speaking, these partition $F$ into two closed sets, and we consider points $x,y$ that minimise the distance between the components. In Theorem 4.6 we show that in any complete metric space, for any compact purely $1$-unrectifiable set $F\subset X$ and $R,r>0$, there exist Besicovitch partitions with $\sum d(x_{i},y_{i})\sim\mathcal{H}^{1}(F)$ such that $d(x_{i},y_{i})<r$ and moreover the balls $B(x_{i},Rd(x_{i},y_{i}))$ (for different $i$) are pairwise disjoint. On the other hand, in Corollary 3.10, we show that if all tangents of a space are connected, then in fact there exists a set of large measure and an $R>0$ such that all tangents to points in this set are connected in a ball of radius $R$. Thus, we again arrive to the desired contradiction.
To conclude the introduction we mention the related work of Preiss and Tišer [12] that gives a different adaptation of Besicovitch’s circle pair argument (to metric spaces) in order to make progress on the Besicovitch one-half conjecture. These results, and the progress towards the conjecture, were recently extended further by De Lellis, Glaudo, Massaccesi and Vittone [6].
The structure of the paper is as follows. After recalling definitions and fixing some notation in §2, we present the definition of a metric tangent and recall various properties from [2]. Theory that is specific to connected tangents is developed in §3.
In §4 we prove our strengthened version of Besicovitch’s circle pair result. Finally, in Theorem 5.1 in §5, we combine all of our ideas and perform the contradiction argument described above. This gives the final piece required to prove Theorem 1.1.
2. Preliminaries
In all the following, let $(X,d)$ be a metric space. For $E\subset X$ and $a\in X$, let
$\textnormal{dist}(a,E):=\inf\{(a,x):x\in E\}.$
Set
$\textnormal{diam}(E):=\sup\{d(x,y):x,y\in E\}.$
We use $B(x,r)$ to denote the closed ball centred at $x$ and of radius $r>0$ and $U(x,r)$ to denote the corresponding open ball. For $\varepsilon>0$, we will denote the $\varepsilon$-neighbourhood of $E$ via an underlined subscript
$E_{\boldsymbol{\underline{\varepsilon}}}:=\{x\in X:\textnormal{dist}(x,E)\leq\varepsilon\}.$
For two sets $A,B\subset X$, we define their Hausdorff distance as
$d_{H}(A,B)=\inf\{\varepsilon>0:A\subset B_{\boldsymbol{\underline{\varepsilon}}}\text{ and }B\subset A_{\boldsymbol{\underline{\varepsilon}}}\}.$
2.1. General Measure Theory
Let $\mathcal{M}(X)$ denote the set of all Borel regular measures on $X$ and $\mathcal{M}_{\text{loc}}(X)\subset\mathcal{M}(X)$ be the set of all those that are finite on all bounded sets.
The support of a measure $\mu\in\mathcal{M}(X)$ is a subset of $X$ and is defined as
$\operatorname{supp}\mu:=\{x\in X:\text{ for all open neighbourhoods }U_{x}\text{ of }x,\,\mu(U_{x})>0\}.$
We may restrict a measure to a set $E\subset X$ defined as $\mu\llcorner E(A):=\mu(E\cap A)$ and pushforward a measure by a map $f$ defined as $f_{\#}\mu(A):=\mu(f^{-1}(A))$.
2.2. Rectifiability
Definition 2.1 (Hausdorff measure).
For $s\geq 0$, $\delta\in[0,\infty]$ and a set $E\subset X$, let
$\mathcal{H}_{\delta}^{s}(E)=\inf\left\{\sum_{i=1}^{\infty}\textnormal{diam}(A_{i})^{s}:\,E\subset\bigcup A_{i},\,\textnormal{diam}(A_{i})<\delta\right\}.$
Then the $s$-dimensional Hausdorff measure of $E$ is defined as
$\mathcal{H}^{s}(E):=\lim_{\delta\rightarrow 0}\mathcal{H}^{s}_{\delta}(E).$
For any $s$, Hausdorff measure is a Borel regular measure [10, Corollary 4.5].
Lemma 2.2.
Let $E\subset X$ be $\mathcal{H}^{s}$-measurable with $\mathcal{H}^{s}(E)<\infty$. Then for any $\varepsilon>0$ we can find a $\delta_{0}>0$ such that for any collection of sets of the form $\{A_{i}\subset X:\textnormal{diam}(A_{i})\leq\delta_{0}\}_{i\in\mathbb{N}}$ satisfies
$\mathcal{H}^{s}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})\leq\sum_{i}\textnormal{diam}(A_{i})^{s}+\varepsilon.$
Proof.
By the definition of Hausdorff measure, $\mathcal{H}^{s}_{\delta}(E)$ is monotonically increasing as $\delta\rightarrow 0$. By assumption, $\mathcal{H}^{s}(E)<\infty$ hence for any $\varepsilon>0$ we can find a $\delta_{0}>0$ such that
$\mathcal{H}^{s}(E)\leq\mathcal{H}^{s}_{\delta}(E)+\varepsilon\qquad\text{for all }\delta\leq\delta_{0}.$
Let $\{A_{i}\subset X:\textnormal{diam}(A_{i})\leq\delta_{0}\}_{i\in\mathbb{N}}$, then
$\mathcal{H}^{s}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})+\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})$ $=\mathcal{H}^{s}(E)$
$\leq\mathcal{H}^{s}_{\delta}(E)+\varepsilon$
$=\mathcal{H}^{s}_{\delta}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})+\mathcal{H}^{s}_{\delta}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})+\varepsilon$
$\leq\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})+\sum_{i}\textnormal{diam}(A_{i})^{s}+\varepsilon.$
Subtracting $\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})$ from both sides (which is finite by assumption) gives the desired result. ∎
Definition 2.3 (Lipschitz Map).
Given a metric space $(X,d)$ and a set $A\subset\mathbb{R}^{n}$, a map $f:A\rightarrow X$ is Lipschitz if there exists $K>0$ such that
$d(f(x),f(y))\leq K|x-y|,\qquad\forall x,\,y\in A.$
Definition 2.4 (Rectifiable).
A set $E\subset X$ is $n$-rectifiable if there are at most countably many Lipschitz maps $f_{i}:A_{i}\rightarrow X$, with $A_{i}\subset\mathbb{R}^{n}$, such that
$\mathcal{H}^{n}(E\setminus\bigcup_{i}f_{i}(A_{i}))=0.$
A set $P\subset X$ is purely $n$-unrectifiable if $\mathcal{H}^{n}(P\cap E)=0$ for all $n$-rectifiable $E\subset X.$
Lemma 2.5.
Let $E\subset X$ be $\mathcal{H}^{n}$-measurable with $\mathcal{H}^{n}(E)<\infty$. There exists a decomposition $E=R\cup P$ where $R$ is $n$-rectifiable and $P$ is purely $n$-unrectifiable.
Proof.
The proof for $X=\mathbb{R}^{m}$ is given in [10, Theorem 15.6]. The proof for general $X$ is identical. ∎
2.3. Density and Doubling
Definition 2.6 (Density).
For $\mu\in\mathcal{M}(X)$, positive number $s\in[0,\infty)$ and point $x\in X$, we can define the upper and lower $s$-dimensional Hausdorff densities of $\mu$ at $x$ as
$\Theta^{*s}(\mu,x)=\limsup_{r\rightarrow 0}\frac{\mu(B(x,r))}{(2r)^{s}};$
$\Theta^{s}_{*}(\mu,x)=\liminf_{r\rightarrow 0}\frac{\mu(E\cap B(x,r))}{(2r)^{s}}.$
If the upper and lower densities agree then the common value is the $s$-dimensional density of $\mu$ at $x$ and is denoted $\Theta^{s}(\mu,x).$ If $E\subset X$ we write $\Theta^{*s}(E,x)$, $\Theta^{s}_{*}(E,x)$ and $\Theta^{s}(E,x)$ for $\Theta^{*s}(\mathcal{H}^{n}|_{E},x)$, $\Theta^{s}_{*}(\mathcal{H}^{n}|_{E},x)$ and $\Theta^{*}(\mathcal{H}^{n}|_{E},x)$ respectively.
Theorem 2.7 ([10] Theorem 6.2).
Let $E\subset X$ be such that $0<\mathcal{H}^{s}(E)<\infty$ then
  1. (i)
    $2^{-s}\leq\Theta^{*s}(E,x)\leq 1$ for $\mathcal{H}^{s}$-almost all $x\in E$;
  2. (ii)
    If $E$ is $\mathcal{H}^{s}$-measurable, then $\Theta^{*s}(E,x)=0$ for $\mathcal{H}^{s}$-almost all $x\in X\backslash E.$
Corollary 2.8.
Let $E,F\subset X$ be $\mathcal{H}^{n}$-measurable with $F\subset E$ and $\mathcal{H}^{n}(E)<\infty$. Then for $\mathcal{H}^{n}$-a.e. $x\in F$,
$\Theta^{*n}(F,x)=\Theta^{*n}(E,x)\qquad\text{and}\qquad\Theta^{n}_{*}(F,x)=\Theta^{n}_{*}(E,x).$
Definition 2.9 (Ahlfors $n$-regular measures).
A metric space $X$ along with a measure $\mu\in\mathcal{M}_{loc}(X)$ is $C$-Ahlfors $n$-regular if for all $r>0$,
$\frac{1}{C}r^{n}\leq\mu(B(x,r))\leq Cr^{n}\qquad\forall\,x\in\operatorname{supp}\mu.$
Denote the set of all $C$-Ahlfors regular metric measure spaces as $\mathcal{A}^{n}(C)$.
Finally we have the notion of doubling for measures.
Definition 2.10 (Doubling).
For $K\geq 1$, $\mu\in\mathcal{M}(X)$ is $K$-doubling if
[…]
or asymptotically doubling if
$\limsup_{r\rightarrow 0}\frac{\mu(B(x,\,2r))}{\mu(B(x,\,r))}<\infty\qquad\textnormal{ for }\mu-a.e.\,x\in X.$
Remark 2.11.
Let $E\subset X$ be $\mathcal{H}^{n}$-measurable with $\mathcal{H}^{n}(E)<\infty$ and $\Theta^{n}_{*}(E,x)>0$ for $\mathcal{H}^{n}$-a.e. $x\in E$. Then $\mathcal{H}^{n}\llcorner E$ is asymptotically doubling. Indeed, By Theorem 2.7(i), for $\mathcal{H}^{n}$-a.e. $x\in E$ we have $\Theta^{*n}(E,x)\leq 1$, so for any such $x$ with $\Theta^{n}_{*}(E,x)>0$ we have:
$\limsup_{r\rightarrow 0}\frac{\mathcal{H}^{n}(E\cap B(x,2r))}{\mathcal{H}^{n}(E\cap B(x,r))}$ $=2^{n}\limsup_{r\rightarrow 0}\frac{\mathcal{H}^{n}(E\cap B(x,2r))}{(2\cdot 2r)^{n}}\cdot\frac{(2r)^{n}}{\mathcal{H}^{n}(E\cap B(x,r))}$
$\leq\frac{2^{n}}{\Theta_{*}^{n}(E,x)}<\infty.$
Theorem 2.12 ([4] Section 3.4).
Let $\mu\in\mathcal{M}(X)$ be asymptotically doubling and $E\subset X$. Then for $\mu$-a.e. $x\in E$
$\lim_{r\rightarrow 0}\frac{\mu(E\cap B(x,r))}{\mu(B(x,r))}=1.$
Any such $x\in S$ is known as a Lebesgue density point of $E$.
2.4. Pointed Gromov-Hausdorff distance
To develop a theory of metric tangents, we require a few standard tools from metric geometry.
Definition 2.13 (Pointed metric (measure) space).
A pointed metric space $(X,\,d,\,x)$ is an ordered triple consisting of a metric space $(X,\,d)$ and a point $x\in X$. A pointed metric measure space $(X,\,d,\,\mu,\,x)$ consists of a complete and separable pointed metric space $(X,x)$ and a measure $\mu\in\mathcal{M}_{loc}(X)$.
An isometric embedding $(X,\,d,\,x)\rightarrow(Y,\,\rho,\,y)$ of pointed metric spaces is an isometric embedding $(X,\,d)\rightarrow(Y,\,\rho)$ such that $x=y.$
Notation 2.14.
When the meaning is clear, we will omit the metric and abbreviate $(X,d,x)$ or $(X,d,\mu,x)$ to $(X,x)$ or $(X,\mu,x)$ respectively.
We first require a pointed notion of distance for sets that share the same ambient space.
Definition 2.15 (Pointed Hausdorff distance).
Let $(Z,d)$ be a metric space and $X,\,Y\subset Z$. Then for $z\in Z$, we define the pointed Hausdorff distance between $X$ and $Y$ as
[…]
We can now define the pointed Gromov-Hausdorff distance which allows us to compare sets across different ambient metric spaces.
Notation 2.16.
In general we shal not distinguish between the original objects $X$ and $\mu$ and their isometric images $\iota(X)\subset Z$ and $\iota_{\#}\mu$.
Definition 2.17 (Pointed Gromov-Hausdorff distance).
Let $(X,\,x)$ and $(Y,\,y)$ be pointed metric spaces. We define the pointed Gromov-Hausdorff distance as
[…]
Finally, let us define the space on which $d_{pGH}$ is a metric. Recall that a metric space is proper if all closed balls are compact.
Corollary 2.18 (Corollary 2.25 [2]).
If $\mathbb{M}_{p}$ is the set of isometry classes of proper pointed metric spaces, then $(\mathbb{M}_{p},\,d_{pGH})$ is a complete and seperable metric space. In particular, if $(X,\,x)$ and $(Y,\,y)$ are proper with
$d_{pGH}((X,x),(Y,y))=0$
then there is an isometry $(X,x)\rightarrow(Y,y)$.
2.5. Convergence of measures
Let $C_{bs}(X)$ denote the space of bounded and continuous maps $f:X\to\mathbb{R}$ with bounded support.
Definition 2.19 (Weak* convergence).
A sequence $\mu_{i}\in\mathcal{M}_{loc}(X)$ weak* converges to $\mu\in\mathcal{M}_{loc}(X)$ if
[…]
for all $f\in C_{bs}(X)$. We use $\xrightharpoonup{*}$ to denote this convergence.
Proposition 2.20 ([7] Proposition 4.26).
Let $\{\mu_{i}\},\mu\in\mathcal{M}_{loc}(X)$. The following are equivalent:
  1. (1)
    $\mu_{i}\xrightharpoonup{*}\mu$.
  2. (2)
    If $x_{i}\rightarrow x\in X$, then for every $r>0$,
    (2.2) $\mu(U(x,r))\leq\liminf_{i\rightarrow\infty}\mu_{i}(U(x_{i},r))$
    and
    (2.3) $\mu(B(x,r))\geq\limsup_{i\rightarrow\infty}\mu_{i}(B(x_{i},r)).$
  3. (3)
    If $B(x,r)$ is such that with $\mu(B(x,r)\backslash U(x,r))=0$, then
    $\mu(B(x,r))=\lim_{i\rightarrow\infty}\mu_{i}(B(x_{i},r)).$
Moreover, if $\mu_{i}\xrightharpoonup{*}\mu$, then for every $x\in\operatorname{supp}\mu$, there exists a sequence $x_{i}\in\operatorname{supp}\mu_{i}$ such that $x_{i}\rightarrow x$.
In general it is not true that if there exists a sequence $\mu_{i}$ with $x_{i}\in\operatorname{supp}\mu_{i}$ such that $\mu_{i}\xrightharpoonup{*}\mu$ and $x_{i}\rightarrow x$ then $x\in\operatorname{supp}\mu$. However, we do have this for Ahlfors regular measures.
Proposition 2.21.
Let $\mu_{i}\in\mathcal{A}^{n}(C)$ in a metric space $X$ be such that $\mu_{i}\xrightharpoonup{*}\mu$. Then $\mu\in\mathcal{A}^{n}(C)$ and if $x_{i}\in\operatorname{supp}\mu_{i}$ is such that $x_{i}\rightarrow x$, then $x\in\operatorname{supp}\mu$.
Proof.
By Proposition 2.20, for any $x\in\operatorname{supp}\mu$ there exists a sequence $x_{i}\in\operatorname{supp}\mu_{i}$ such that $x_{i}\rightarrow x$. Thus for any $\varepsilon>0$ we have
$\mu(B(x,r))\leq\mu(U(x,r+\varepsilon))\leq\liminf_{i\rightarrow\infty}\mu_{i}(B(x_{i},r+\varepsilon))\leq C(r+\varepsilon)^{n}$
and
$\mu(B(x,r))\geq\limsup_{i\rightarrow\infty}\mu_{i}(B(x_{i},r))\geq r^{n}/C.$
Letting $\varepsilon\rightarrow 0$ demonstrates that $\mu\in\mathcal{A}^{n}(C)$.
For the second claim, fix $r>0$ and let $i_{0}>0$ be sufficiently large that $B(x_{i},r)\subset B(x,2r)$ for all $i\geq i_{0}$. Thus
$\mu(B(x,2r))\geq\limsup_{i\rightarrow\infty}\mu_{i}(B(x,2r))\geq\limsup_{i\rightarrow\infty}\mu_{i}(B(x_{i},r))\geq r^{n}/C.$
By the arbitrariness of $r$, we have demonstrated that $x\in\operatorname{supp}\mu$. ∎
As previously intimated, we will be working in terms of tangent measures as opposed to sets.
Definition 2.22.
Let $x\in X$ and $\mu,\nu\in\mathcal{M}(X)$. For $L,\,r>0$ define
$F^{L,\,r}_{x}(\mu,\nu):=\sup\left\{\int f\,d(\mu-\nu):f\in\textnormal{Lip}_{L}(X;[-1,1])\,\textnormal{ with }\operatorname{supp}f\subset B(x,\,r)\right\}.$
Further, where the infimum exists let
$F_{x}(\mu,\nu):=\inf\left\{\varepsilon\in(0,1/2):F_{x}^{1/\varepsilon,1/\varepsilon}(\mu,\nu)<\varepsilon\right\}$
and if no such infimum exists, set $F_{x}(\mu,\nu)=1/2$.
We have the following useful equivalence:
Proposition 2.23 (Proposition 2.13 [2]).
Convergence in $F_{x}$ implies weak* convergence. If $X$ is complete and separable then the converse is also true.
We now introduce a measured version of $d_{pGH}$.
Definition 2.24.
For pointed metric measure spaces $(X,\mu,z)$ and $(Y,\nu,y)$, define
[…]
This is a complete and separable metric on the set of isometry classes of pointed metric measure spaces $\mathbb{M}_{pm}$, see [2, Corollary 4.6]. We are now ready to introduce the metric that will be used to define tangent measures in the next section.
Definition 2.25.
Let $(X,\mu,z)$ and $(Y,\nu,y)$ be pointed metric measure spaces and define
[…]
Let $\mathbb{M}_{*}$ be the set of pointed metric measure spaces under the equivalence relation
$(X,\mu,x)\sim(Y,\nu,y)\text{ if }(\operatorname{supp}\mu\cup\{x\},\mu,x)\textnormal{ is isometric to }(\operatorname{supp}\nu\cup\{y\},\nu,y).$
Then as shown in [2, Corollary 4.13], $(\mathbb{M}_{*},d_{*})$ is a complete and separable metric space. In light of this, to simplify notation we will sometimes write $(\mu,x)$ in place of $(X,\mu,x)$.
The final piece of the puzzle will be to extract some relationship between the convergence of measures and the convergence of their supports. Although $d_{*}$ and $d_{pmGH}$ are not equivalent, we do have a strong relationship as outlined in the following key result.
Theorem 2.26 (Theorem 4.11 [2]).
For $\varepsilon>0$, let $(X,\mu,x),(Y,\nu,y)\in\mathbb{M}_{*}$ be pointed metric measure spaces with $x\in\operatorname{supp}\mu$, $y\in\operatorname{supp}\nu$ and
$d_{*}((X,\mu,x),(Y,\nu,y))<\varepsilon.$
There exists compact $x\in K_{\mu}\subset\operatorname{supp}\mu\cap B(x,1/\varepsilon)$ and $y\in K_{\nu}\subset\operatorname{supp}\nu\cap B(y,1/\varepsilon)$ such that
  1. M1.
    $\max\{\mu(B(x,1/\varepsilon)\backslash K_{\mu}),\nu(B(y,1/\varepsilon)\backslash K_{\nu})\}<\varepsilon$;
  2. M2.
    $d_{pmGH}((K_{\mu},\mu\llcorner K_{\mu},x),(K_{\nu},\nu\llcorner K_{\nu},y))<3\varepsilon$.
Conversely, if there exist compact $K_{\mu},\,K_{\nu}$ as above satisfying M1. with
$d_{pmGH}((K_{\mu},\mu\llcorner K_{\mu},x),(K_{\nu},\nu\llcorner K_{\nu},y))<\varepsilon$
then
$d_{*}((X,\mu,x),(Y,\nu,y))<3\varepsilon.$
That is, $d_{*}$ is bi-Lipschitz equivalent to the metric defined by taking the infimum over all $\varepsilon\in(0,1/2)$ such that M1. and M2. hold.
We conclude this section with the following proposition that allows us to work entirely in some common space $(Z,z)$.
Proposition 2.27 (Proposition 2.20 [2]).
Let $\{(X_{i},\mu_{i},x_{i})\}_{i}$ be a sequence of pointed metric measure spaces. We have that
$d_{*}((X_{i},\mu_{i},x_{i}),(X,\mu,x))\rightarrow 0,\quad\text{respectively}\quad d_{pmGH}((X_{i},\mu_{i},x_{i}),(X,\mu,x))\rightarrow 0,$
for some pointed metric measure space $(X,\mu,x)$ if and only if there exists a complete and separable pointed metric space $(Z,z)$ and isometric embeddings $(X_{i},x_{i})\rightarrow(Z,z)$ and $(X,x)\rightarrow(Z,z)$ such that
$F_{z}(\mu_{i},\mu)\rightarrow 0,\qquad\text{or respectively}\qquad d_{H(z)}(X,Y)+F_{z}(\mu,\,\nu)\rightarrow 0.$
2.6. Tangents of metric measure spaces
Tangent measures were first introduced by Preiss [11] in Euclidean space as a natural setting in which to explore questions around density. We refer the reader to [10, Page 184] for an introduction to tangent measures. With the notion of convergence of metric measure spaces established in the previous section, the definition of tangent measures in metric spaces follows analogously. First, let $(\mu,d,x)\in\mathbb{M}_{*}$ with $x\in\operatorname{supp}\mu$ and for $r>0$, define
$T_{r}(\mu,d,x):=\left(\frac{\mu}{\mu(B(x,r))},\frac{d}{r},x\right).$
Definition 2.28 (Tangent measure).
For $(\mu,d,x)\in\mathbb{M}_{*}$, we say that $(\nu,\rho,y)\in\mathbb{M}_{*}$ is a tangent measure if there exists a sequence $\{r_{i}\}$ with $r_{i}\rightarrow 0$ such that
$d_{*}(T_{r_{i}}(\mu,d,x),(\nu,\rho,y))\rightarrow 0\qquad\text{as }i\rightarrow\infty.$
The set of all tangent measures for $(\mu,d,x)$ is denoted $\textnormal{Tan}(\mu,d,x)$.
Corollary 2.29 (Corollary 4.7 [2]).
Let $(\mu,d,x)\in\mathbb{M}_{*}$ and let $S\subset\operatorname{supp}\mu$ be $\mu$-measurable with $x$ a Lebesgue density point of $S$. Then
$\textnormal{Tan}(\mu\llcorner S,x)=\textnormal{Tan}(\mu,x).$
We also have the following compactness result.
Proposition 2.30 (Proposition 5.3 [2]).
Let $X$ be a metric space and let $\mu\in\mathcal{M}_{loc}(X)$ be asymptotically doubling. For $\mu$-a.e. $x\in\operatorname{supp}\mu$ and every $r_{0}>0,$
[…]
In particular, for any $x\in\operatorname{supp}\mu$ satisfying the above, $\textnormal{Tan}(\mu,x)$ is a non-empty compact metric space when equipped with $d_{*}$ and
$\forall\,\delta>0,\exists\,r_{x}>0:d_{*}(T_{r}(\mu,x),\textnormal{Tan}(\mu,x))\leq\delta\,\forall\,r\in(0,r_{x}).$
Due to the rescaling of the metric in the definition and the nature of embeddings we can derive the fact that the tangent space is closed under rescaling of metrics and translation of points.
Lemma 2.31.
Let $(X,d,\mu)$ be a metric measure space such that $\mu$ is asymptotically doubling.
  1. (1)
    For $\mu$-a.e. $x\in\operatorname{supp}\mu$ and $s>0$, there exists $C>0$ such that
    $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)\quad\implies\quad(C\nu,s\rho,y)\in\textnormal{Tan}(\mu,x)$
  2. (2)
    For $\mu$-a.e. $x\in\operatorname{supp}\mu$,
    $(\nu,y)\in\textnormal{Tan}(\mu,x)\textnormal{ and }z\in\operatorname{supp}\nu\quad\implies\quad(\nu,z)\in\textnormal{Tan}(\mu,x).$
Proof.
Proof of (2) may be found in [2, Proposition 5.5] so we just present a proof of (1). If $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)$ then there exists $r_{i}\rightarrow 0$ such that
$d_{*}(T_{r_{i}}(\mu,d,x),(\nu,\rho,y))\rightarrow 0.$
By Proposition 2.27 there exists a complete and separable metric space $(Z,\lambda,z)$ and isometries
$\left(\operatorname{supp}\frac{\mu}{\mu(B_{d}(x,r_{i}))},\frac{d}{r_{i}},x\right),(\operatorname{supp}\nu,\rho,y)\rightarrow(Z,\lambda,z)$
such that
$F_{z}\left(\frac{\mu}{\mu(B_{d}(x,r_{i}))},\nu\right)\rightarrow 0.$
Let
$C_{i}=\frac{\mu(B_{d}(x,r_{i}))}{\mu(B_{sd}(x,r_{i}))}.$
Since $\mu$ is asymptotically doubling we can set (up to subsequence)
$C=\lim_{i\to\infty}C_{i}>0.$
Consider now the sequence $r_{i}/s\rightarrow 0$ and note that
$T_{r_{i}/s}(\mu,d,x)$ $=\left(\frac{\mu}{\mu(B_{d}(x,r_{i}/s))},\frac{sd}{r_{i}},x\right)$
$=\left(\frac{\mu}{\mu(B_{sd}(x,r_{i}))},\frac{sd}{r_{i}},x\right)$
$=T_{r_{i}}(\mu,sd,x).$
Here the subscript denotes the metric that defines the ball. By pre and post-scaling the previous isometries by $s$, we can construct new isometries
$\left(\operatorname{supp}\frac{\mu}{\mu(B_{sd}(x,r_{i}))},\frac{sd}{r_{i}},x\right),(\operatorname{supp}\nu,s\rho,y)\rightarrow(Z,s\lambda,z).$
We will write $\mu^{s}$ and $\nu^{s}$ for the pushforward of $\mu,\,\nu$ under these isometries in $(Z,s\lambda)$ and scaling $\nu^{s}$ we have,
$d_{*}\left(T_{r_{i}}(\mu,sd,x),\left(C\nu,s\rho,y\right)\right)$ $\leq F_{z}\left(\frac{\mu^{s}}{\mu(B_{sd}(x,r_{i}))},C\nu^{s}\right)$
$\leq F_{z}\left(\frac{\mu^{s}}{\mu(B_{sd}(x,r_{i}))},C_{i}\nu^{s}\right)+F_{z}\left(C_{i}\nu^{s},C\nu^{s}\right)$
$\leq F_{z}\left(\frac{\mu^{s}}{\mu(B_{sd}(x,r_{i}))},C_{i}\nu^{s}\right)+|C-C_{i}|.$
Recall that by Proposition 2.23 we have that convergence in $F_{z}$ is equivalent to weak* convergence. Thus for any $f\in C_{bs}(Z)$,
$\int f(z)\,d\,\left(\frac{\mu^{s}}{\mu(B_{sd}(x,r_{i}))}-C_{i}\nu^{s}\right)(z)=C_{i}\int f(sz)\,d\,\left(\frac{\mu}{\mu(B_{d}(x,r_{i}))}-\nu\right)(z)\rightarrow 0$
which combined with $|C-C_{i}|\to 0$ concludes the proof. ∎
The upper and lower density of the measure at a point has implications on the regularity of tangents at that point.
Lemma 2.32.
For $(\mu,x)\in\mathbb{M}_{*}$ with $0<\Theta^{s}_{*}(\mu,x),\,\Theta^{s*}(\mu,x)<\infty$, we have that
$\frac{\Theta^{s}_{*}(\mu,x)}{\Theta^{s*}(\mu,x)}r^{s}\leq\nu(B(y,r))\leq\frac{\Theta^{s*}(\mu,x)}{\Theta^{s}_{*}(\mu,x)}r^{s}\qquad\forall(\nu,y)\in\textnormal{Tan}(\mu,x).$
In particular, if $0<\Theta^{s}_{*}(\mu,x),\,\Theta^{s*}(\mu,x)<\infty$ for $\mu$-a.e. $x\in\operatorname{supp}\mu$, then for $\mu$-a.e. $x\in\operatorname{supp}\mu$,
$\textnormal{Tan}(\mu,x)\subset\mathcal{A}\left(\frac{\Theta^{s*}(\mu,x)}{\Theta^{s}_{*}(\mu,x)}\right).$
Proof.
By Proposition 2.27 and Proposition 2.23, we have that for any $(\nu,y)\in\textnormal{Tan}(\mu,x)$ there exists a sequence $r_{i}\rightarrow 0$, a metric space $(Z,\lambda,z)$ and isometries $(T_{r_{i}}(\mu,d,x)),(\nu,\rho,y)\rightarrow(Z,\lambda,z)$ such that
$\frac{\mu_{i}}{\mu(B(x,r_{i}))}\xrightharpoonup{*}\nu.$
Thus by (2.2) we have
$\nu(B(y,R))\leq\nu(U(y,R+\varepsilon))$ $\leq\liminf_{i\rightarrow\infty}\frac{\mu(B(x,(R+\varepsilon)r_{i}))}{\mu(B(x,r_{i}))}$
$=(R+\varepsilon)^{s}\liminf_{i\rightarrow\infty}\frac{\mu(B(x,(R+\varepsilon)r_{i}))}{(2(R+\varepsilon)r_{i})^{s}}\cdot\frac{(2r_{i})^{s}}{\mu(B(x,r_{i}))}$
$\leq(R+\varepsilon)^{s}\frac{\Theta^{s*}(\mu,x)}{\Theta^{s}_{*}(\mu,x)}\xrightarrow{\varepsilon\rightarrow 0}R^{s}\frac{\Theta^{s*}(\mu,x)}{\Theta^{s}_{*}(\mu,x)}.$
Similarly, by (2.3),
$\nu(B(y,R))\geq\limsup_{i\rightarrow\infty}\frac{\mu(B(x,(R+\varepsilon)r_{i}))}{\mu(B(x,r_{i}))}\geq R^{s}\frac{\Theta_{*}^{s}(\mu,x)}{\Theta^{s*}(\mu,x)}$
which concludes the first statement. The particular case follows from Lemma 2.31(2). ∎
Corollary 2.33.
For an $\mathcal{H}^{s}$-measurable set $E\subset X$ with $\mathcal{H}^{s}(E)<\infty$, we can combine Lemma 2.32 and Theorem 2.7, to see that for $\mathcal{H}^{s}$-a.e. $x\in E$,
$R^{s}\Theta_{*}^{s}(\mathcal{H}^{s}\llcorner E,x)\leq\nu(B(y,R))\leq R^{s}\frac{1}{\Theta^{s}_{*}(\mathcal{H}^{s}\llcorner E,x)}$
for all $(\nu,y)\in\textnormal{Tan}(\mathcal{H}^{s}\llcorner E,x)$.
3. Connected Tangents of Metric Spaces
In this section we show how the hypothesis that all tangent spaces are connected improves to a stronger, quantitative notion of connectedness.
Let $\mathcal{C}$ be the collection of connected separable metric spaces and let $\mathcal{C}^{*}=\{(\mu,d,x)\in\mathbb{M}_{*}:\operatorname{supp}\mu\in\mathcal{C}\}$ equipped with $d_{*}$.
Definition 3.1 (Quasi-path-connected).
Let $\delta>0$. We say that a metric space $(X,d)$ admits a $\delta$-quasi-path between $a,\,b\in X$ if there exists a finite collection of points $\{x_{n}\}_{n=1}^{N}\subset X$ such that $x_{1}=a$, $x_{N}=b$ and $d(x_{n},\,x_{n+1})\leq\delta d(a,b)$ for each $n$. We have the following refinements:
  1. (1)
    $\mathcal{Q}(\delta,R)$ is the collection of metric spaces $(X,d)$ such that for all $a,b\in X$, $B(a,Rd(a,b))$ admits a $\delta$-quasi-path between $a$ and $b$. We say that such an $X$ is $R$-locally $\delta$-quasi-path-connected.
  2. (2)
    $\mathcal{Q}_{p}(\delta,R)$ is the collection of pointed metric spaces $(X,d,x)$ such that for all $y\in X$, $B(x,Rd(x,y))$ admits a $\delta$-quasi-path between $x$ and $y$.
  3. (3)
    $\mathcal{Q}_{*}(\delta,R)=\{(\mu,y)\in\mathbb{M}_{*}:(\operatorname{supp}\mu,y)\in\mathcal{Q}_{p}(\delta,R)\}.$
  4. (4)
    $\mathcal{Q}_{*}(\delta,R,C)=\mathcal{Q}_{*}(\delta,R)\cap\mathcal{A}(C).$
Lemma 3.2.
For $\delta,R,C>0$, let $(X,d,\mu,a)\notin\mathcal{Q}_{*}(\delta,\,R,\,C)$. Then there exists a point $b\in\operatorname{supp}\mu$ such that there exists a decomposition of $\operatorname{supp}\mu\cap B(a,Rd(a,b))$ into sets $A$ and $B$ containing $a$ and $b$ respectively with
$\textnormal{dist}(A,B)>\delta d(a,b).$
Proof.
Suppose $(X,d,\mu,a)\notin\mathcal{Q}_{*}(\delta,\,R)$. Then there exists a point $b\in\operatorname{supp}\mu$ such that there is no $\delta$-quasi-path between $a$ and $b$ in $B(a,Rd(a,b))$. Define
$A=\{x\in\operatorname{supp}\mu\cap B(a,Rd(a,b)):\\ \textnormal{there exists a $\delta$-quasi-path between $a$ and $x$ in $B(a,Rd(a,b))$}\}$
and $B=\operatorname{supp}\mu\cap B(a,Rd(a,b))\setminus A$. Then $a\in A$ and $b\in B$.
Suppose that $\textnormal{dist}(A,B)\leq\delta(a,b)$. Then, since $\mu\in\mathcal{A}(C)$, $A$ and $B$ are compact. Thus there exists $a^{\prime}\in A$ and $b^{\prime}\in B$ such that $d(a^{\prime},b^{\prime})\leq\delta d(a,b)$. Then there exists a $\delta$-quasi-path between $a$ and $b$ in $B(a,Rd(a,b))$, a contradiction. ∎
Lemma 3.3.
Suppose $E\subset X$ is such that there exists a $\delta$-quasi-path connecting $x,y\in E$ in $B(x,Rd(x,y))$ for some $R>1$. Suppose also that $F\subset X$ satisfies
(3.1) $d_{H(x)}(E,F)<\varepsilon d(x,y),\qquad\text{where }\varepsilon<\min\left\{\frac{1}{(R+1)d(x,y)^{2}},1\right\}.$
Then $F\cup\{x,y\}$ contains a $(\delta+2\varepsilon)$-quasi-path connecting $x$ and $y$ in
$B(x,(R+\varepsilon)d(x,y)).$
Proof.
Let $\{x_{n}\}\in E$ be the $\delta$-quasi-path connecting $x,y\in E$. Note that
$\frac{1}{\varepsilon d(x,y)}>(R+\varepsilon)d(x,y).$
Thus by (3.1), each $x_{n}$ is contained in $B(x,1/\varepsilon d(x,y))$ and by (3.1) there exists an $\hat{x}_{n}\in F$ such that
$d(x_{n},\hat{x}_{n})<\varepsilon d(x,y).$
Thus, the triangle inequality gives us
$d(\hat{x}_{n},\hat{x}_{n+1})\leq d(\hat{x}_{n},x_{n})+d(x_{n},x_{n+1})+d(\hat{x}_{n+1},x_{n+1})\leq(\delta+2\varepsilon)d(x,y).$
Further, by assumption we have
$d(x,x_{n})\leq Rd(x,y)\qquad\forall n\in\mathbb{N}$
thus another application of the triangle inequality tells us that $\{\hat{x}_{n}\}\cup\{x,y\}$ is the required $(\delta+2\varepsilon)$-quasi-path in $(F\cup\{x,y\})\cap B(x,(R+\varepsilon)d(x,y))$. ∎
Lemma 3.4.
Let $(X,\mu,x)$ be a pointed metric space and $\varepsilon,\delta\in(0,1)$. Let $\hat{x}\in B(x,(1/\varepsilon-\delta))$ be such that
(3.2) $\mu(B(\hat{x},\delta))>\varepsilon.$
If $K_{\mu}\subset X$ is such that
(3.3) $\mu\left(B\left(x,\frac{1}{\varepsilon}\right)\backslash K_{\mu}\right)<\varepsilon,$
then $K_{\mu}$ must contain a point $k_{\mu}\in K_{\mu}$ such that
$d(\hat{x},k_{\mu})<\delta.$
Proof.
Because $B(\hat{x},\delta)\subset B(x,1/\varepsilon)$, combine (3.2) with (3.3) to see that there must be some point $k_{\mu}\in K_{\mu}\cap B(\hat{x},\delta)$. ∎
Corollary 3.5.
For $R>1$, let $(X,\mu,x)$ be a metric measure space, $y\in\operatorname{supp}\mu$ and suppose that $\operatorname{supp}\mu$ contains a $\delta$-quasi-path $\{x_{n}\}$ in $B(x,Rd(x,y))$ connecting $x$ and $y$. Suppose also that for $\varepsilon,\delta\in(0,1)$,
(3.4) $\mu(B(x_{n},\delta d(x,y)))>\varepsilon\qquad\forall\,n\in\mathbb{N}.$
Then for any $K_{\mu}\subset X$ containing $x,\,y$ and satisfying
(3.5) $\mu(B(y,(R+1)d(x,y))\backslash K_{\mu})<\varepsilon,$
$K_{\mu}\cap B(y,(R+\delta)d(x,y))$ contains a $3\delta$-quasi-path connecting $y$ and $t$.
Proof.
Lemma 3.4 implies that, for each $n\in\mathbb{N}$, there exists some point $q_{n}\in K_{\mu}\cap B(x_{n},\delta d(x,y))$. Hence,
$d(q_{n},q_{n+1})\leq d(q_{n},x_{n})+d(x_{n},x_{n+1})+d(q_{n+1},x_{n+1})\leq 3\delta d(x,y).$
Thus the collection $\{y,t,q_{n}\}$ forms the required $3\delta$-quasi-path in $K_{\nu}\cap B(y,(R+\delta)d(y,t))$. ∎
In general, $\mathcal{Q}(\delta,R)\subsetneq\mathcal{Q}_{p}(\delta,R)$. Indeed, if $X=\{0\}\cup\{1/2^{n}\}_{n\in\mathbb{N}}$ with the Euclidean metric, then $(X,\|\cdot\|_{2},0)\in\mathcal{Q}_{p}(1,1/2)$ but for any $n$, there is no 1/2-quasi-path from $1/2^{n}$ to $1/2^{n+1}$ so $X\notin\mathcal{Q}(1,1/2)$. However, in the context of tangents, Lemma 2.31(2) shows that $\mathcal{Q}_{p}(\delta,R)$ and $\mathcal{Q}(\delta,R)$ are essentially equivalent.
Corollary 3.6.
Let $(X,d,\mu)$ be an asymptotically doubling metric measure space. Suppose that for $\mu$-a.e. $x\in\operatorname{supp}\mu$ there exists $R_{x},\delta_{x}>0$ such that
$\textnormal{Tan}(\mu,x)\subset\mathcal{Q}_{*}(\delta_{x},R_{x}).$
Then for $\mu$-a.e. $x\in\operatorname{supp}\mu$,
$(\nu,y)\in\textnormal{Tan}(\mu,x)\quad\implies\quad(\operatorname{supp}\nu,y)\in\mathcal{Q}(\delta_{x},R_{x}).$
We now show that if all tangents of a space are connected, then all tangents are in fact quasi-path-connected.
Proposition 3.7.
Let $(X,\,d)$ be a complete metric space and $\mu\in\mathcal{M}_{loc}(X)$. Suppose that for $\mu$-a.e. $x\in\operatorname{supp}\mu$ we have
  1. (i)
    $0<\Theta^{1}_{*}(\mu,x),\,\Theta^{1*}(\mu,x)<\infty$;
  2. (ii)
    $\textnormal{Tan}(\mu,x)\subset\mathcal{C}^{*}.$
Then for $\mu$-a.e. $x\in\operatorname{supp}\mu$ and any $\delta>0$ there exists $R_{x},C_{x}>0$ such that $\textnormal{Tan}(\mu,x)\subset\mathcal{Q}_{*}(\delta,R_{x},C_{x})$.
Proof.
Fix $\delta>0$ and an $x$ that satisfies (i), (ii), has a compact tangent space as given by Proposition 2.30 and such that we may apply Lemma 2.31(1) and (2). Note that $\mu$-a.e. $x\in X$ satisfies theses requirements. First, recall that Lemma 2.32 gives us a $C_{x}>1$ such that
(3.6) $\frac{1}{C_{x}}r\leq\nu(B(z,r))\leq C_{x}r\qquad\forall\,z\in\operatorname{supp}\nu,\,r>0.$
To demonstrate the existence of $R_{x}$ we work by contradiction, see Figure 1 for a geometric description of the proof. If there exists no such $R_{x}>0$, then for any $n\in\mathbb{N}$ there exists a tangent measure $(\nu_{n},\rho_{n},a_{n})\in\textnormal{Tan}(\mu,x)$ such that $(\nu_{n},\rho_{n},a_{n})\notin\mathcal{Q}_{*}(\delta,n,C_{x})$. Thus, by Lemma 3.2, there exists a point $b_{n}\in\operatorname{supp}\nu_{n}$ such that $\operatorname{supp}\nu_{n}\cap B(a_{n},n\rho_{n}(a_{n},b_{n}))$ can be decomposed into sets $A_{n}$ and $B_{n}$ containing $a_{n}$ and $b_{n}$ respectively with
$\textnormal{dist}_{\rho_{n}}(A_{n},B_{n})>\delta\rho_{n}(a_{n},b_{n}).$
We use the $(\nu_{n},a_{n})$ to construct another sequence of elements of $\textnormal{Tan}(\mu,x)$.
By Lemma 2.31(1), we may rescale the metrics of each tangent and suppose that $\rho_{n}(a_{n},b_{n})=1$ (we omit the subsequent scaling of the measure by $C$ for ease of notation). By compactness of $\textnormal{Tan}(\mu,x)$, there exists a subsequence $(\nu_{n_{k}},\rho_{n_{k}},a_{n_{k}})$ such that, for any $\varepsilon>0$, there exists a $K>0$ such that
$d_{*}((\nu_{n_{k}},a_{n_{k}}),(\nu_{n_{l}},a_{n_{l}}))<\varepsilon\qquad\forall\,k,l>K.$
By Lemma 2.27, we may consider $(\nu_{n_{k}},a_{n_{k}})$ to be in a common ambient space $(Z,d_{Z})$ with each $a_{n_{k}}$ mapped to a common point $a\in Z$. For brevity we relabel the subsequence to be indexed once again by $n$.
Finally, by Lemma 2.31(2) we observe that $(\nu_{n},b_{n})$ are also tangent measures. Hence there exists a sequence $\{(\nu_{n},\rho_{n},b_{n})\}_{n\in\mathbb{N}}\subset\textnormal{Tan}(\mu,x)$ such that $\operatorname{supp}\nu_{n}\cap B(a,n)$ has decomposition $A_{n}\cup B_{n}$ satisfying
$\textnormal{dist}_{Z}(A_{n},B_{n})>\delta.$
By compactness of $\textnormal{Tan}(\mu,x)$, $\{(\nu_{n},b_{n})\}$ converges in $d_{*}$ (up to subsequence) to a limit $(\nu,b)\in\textnormal{Tan}(\mu,x)$ with $a\in\operatorname{supp}\nu$. Since $\rho_{n}(a_{n},b_{n})=1$ for each $n$, $d_{Z}(a,b)=1$.
By (ii) we have that $(\nu,b)\in\mathcal{C}^{*}$, thus there must exist a finite $\delta/3$-quasi-path $\{z^{i}\}\subset\operatorname{supp}\nu$ between $a$ and $b$ and a radius $R$ such that $\{z^{i}\}\subset B(b,R)$.
[…]
$d_{Z}(z^{i}_{N},z^{i+1}_{N})\leq d_{Z}(z^{i}_{N},z^{i})+d_{Z}(z^{i},z^{i+1})+d_{Z}(z^{i+1},z^{i+1}_{N})\leq\delta.$
By another application of the triangle inequality we also have that $\{z^{i}_{N}\}\subset B(b,R+\delta)\subset B(b,N)$. This contradicts our initial assumption that no such $\delta$-path exists in $\operatorname{supp}\nu_{N}\cap B(b,N)$.
$R$
$\leq\delta$
$\leq\frac{1}{R}$
$\geq\delta$
$n$
$a$
$b$
$a$
$b$
$b_{n}$
$A_{n}$
$B_{n}$
Figure 1. The contradiction in the proof of Lemma 3.7. On the left, the limit of the sequence $(\nu_{n},b_{n})$ must have a radius which contains a $\delta$-quasi-path. On the right, for sufficiently large $n$, the existence of the $\delta$-quasi-path implies there must be points of $\operatorname{supp}\nu_{n}$ in the gap.
We now decompose $X$ into countably many Borel pieces on which we have a uniform bound on the $R_{x}$ and $C_{x}$ obtained from Proposition 3.7.
Lemma 3.8 (Lemma 5.4 [2]).
Let $(X,d,\mu)$ be a metric measure space and suppose that $\mathcal{S}\subset\mathbb{M}_{*}$. Then for any $k,\varepsilon>0$,
$C_{k,\varepsilon}:=\{x\in\operatorname{supp}\mu:d_{*}(T_{r}(\mu,d,x),\mathcal{S})\leq\varepsilon,\,\forall\,r<k\}$
is a closed subset of $X$. In particular, if $E\subset X$ is a Borel set of points satisfying Proposition 2.30 and $\mathcal{S}\subset\mathbb{M}_{*}$ is closed, then
$\{x\in S:\textnormal{Tan}(\mu,d,x)\subset\mathcal{S}\}$
is Borel.
Lemma 3.9.
For any $\delta<1/2$ and $R,C\geq 1$, $\mathcal{Q}_{*}(\delta,R,C)$ is a closed subset of $(\mathbb{M}_{*},d_{*})$.
Proof.
Take a sequence $(\mu_{i},x_{i})\in\mathcal{Q}_{*}(\delta,R,C)$ that converges to $(\mu,x)\in\mathbb{M}_{*}$. By Proposition 2.27 we may consider the sequence and limit to be contained in complete and separable metric space $(X,d)$ with each $x_{i}$ mapped to $x$. Then by Proposition 2.23 we have $\mu_{i}\xrightharpoonup{*}\mu$.
Let $y\in\operatorname{supp}\mu$. We will demonstrate that there exists a $\delta$-quasi-path between $x$ and $y$.
First, by Proposition 2.21, we know that $\mu\in\mathcal{A}(C)$. By Proposition 2.20, $y\in\operatorname{supp}\mu$ implies that there exists a sequence $y_{i}\in\operatorname{supp}\mu_{i}$ such that $y_{i}\rightarrow y$. Since for each $i$ we have that $\mu_{i}\in\mathcal{Q}_{*}(\delta,R,C)$, we also have a sequence of $\delta$-quasi-paths $\{y_{i}^{n}\}_{n\in\mathbb{N}}\subset\operatorname{supp}\mu_{i}\cap B(x,Rd(x,y_{i}))$.
Let us assume that $\mu\notin\mathcal{Q}_{*}(\delta,R,C)$, then, as in Lemma 3.2, there exists a point $y\in\operatorname{supp}\mu$ and two open sets $A_{1},A_{2}\subset B(x,Rd(x,y))$ such that
  1. (1)
    $\operatorname{supp}\mu\cap B(x,Rd(x,y))\subset A_{1}\cup A_{2}$;
  2. (2)
    $x\in A_{1}$ and $y\in A_{2}$;
  3. (3)
    $\textnormal{dist}(A_{1},A_{2})>\delta d(x,y)$.
Then $B=B(x,Rd(x,y))\backslash A_{1}\cup A_{2}$ is disjoint from $\operatorname{supp}\mu$. Now, because each $\mu_{i}\in\mathcal{Q}_{*}(\delta,R,C)$, no such similar decomposition of $B(x,Rd(x,y))$ is possible, i.e. for each $\mu_{i}$ there exists some $z_{i}\in B\cap\operatorname{supp}\mu_{i}$. Hence, by using the Ahlfors regularity of $\mu_{i}$ we see that we must have
$\mu_{i}(B)>0.$
Thus
$\mu(B)\geq\limsup_{i\rightarrow\infty}\mu_{i}(B)>0$
which directly contradicts item (1), concluding the proof.
Corollary 3.10.
Let $(X,d)$ be a complete metric space and $E\subset X$ be $\mathcal{H}^{1}$-measurable with $\mathcal{H}^{1}(E)<\infty$. Assume that for $\mathcal{H}^{1}$-a.e. $x\in E$,
$\Theta_{*}(E,x)>0,\qquad\text{and}\qquad\textnormal{Tan}(X,d,\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{C}^{*}.$
Then for any $\delta>0$, there exists a compact set $F\subset E$ with $\mathcal{H}^{1}(F)>0$ satisfying the following:
  1. S1.
    There exists $\hat{d}>0$ such that for $\mathcal{H}^{1}$-a.e. $x\in F$,
    $\hat{d}\leq\Theta_{*}(F,x),\qquad\Theta^{*}(F,x)\leq 1.$
  2. S2.
    There exists $R>0$ such that for $\mathcal{H}^{1}$-a.e. $x\in F$,
    $\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,x)\subset\mathcal{Q}_{*}(\delta,R,1/\hat{d})\cap\mathcal{C}^{*}.$
Proof.
Fix $\delta>0$. By Theorem 2.7, we can find a set $E^{\prime}\subset E$ of positive measure with constants $\hat{d},\,r_{0}>0$ such that for all $x\in E^{\prime}$,
$2\hat{d}r\leq\mathcal{H}^{1}(E\cap B(x,r))\leq 4r\qquad\forall\,r<r_{0}.$
For any $R>0$, define the set
$F_{R}=\{x\in E:\textnormal{Tan}(\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{Q}_{*}(\delta,R,1/\hat{d})\}.$
By Proposition 3.7 and Proposition 2.32, we know that $F_{R}$ monotonically increase with $R$ to cover $\mathcal{H}^{1}$-almost all of $E^{\prime}$. By Lemma 3.9 and Lemma 3.8, $F_{R}$ are measurable thus there exists $R>0$ such that
$\mathcal{H}^{1}(E^{\prime}\cap F_{R})>0.$
Because $E^{\prime}\cap F_{R}$ is $\mathcal{H}^{1}$-measurable, we may find some compact subset $F\subset E^{\prime}\cap F_{R}$ with $\mathcal{H}^{1}(F)>0$. Using Corollary 2.8 we have that for $\mathcal{H}^{1}$-a.e. $x\in F$
$\hat{d}\leq\Theta_{*}(F,x),\qquad\Theta^{*}(F,x)\leq 1.$
So $F$ satisfies S1.. Recalling Remark 2.11, $\mathcal{H}^{1}\llcorner E$ is asymptotically doubling, thus by Lemma 2.29, for $\mathcal{H}^{1}$-a.e. $x\in F$,
$\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,x)=\textnormal{Tan}(\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{Q}_{*}(\delta,R,1/\hat{d})\cap\mathcal{C}^{*},$
which gives us S2. as required.
4. Besicovitch Partitions
In this section we show that a large proportion of a purely 1-unrectifiable set in a metric space can be covered by Besicovitch partitions. A Besicovitch partition is an adaptation of Besicovitch’s ideas on circle pairs in order to work with quasi-connected tangents. Indeed, in the next section we will see that a set with a suitable collection of Besicovitch partitions cannot have quasi-path connected tangents.
In place of the Besicovitch covering theorem, we use the Alhfors type estimate (4.1) to bound the size of the cover.
Notation 4.1.
For a collection $\mathcal{A}$ and a set $A\in\mathcal{A}$, let
$A_{\cap}:=\{B\in\mathcal{A}:A\cap B\neq\emptyset\}.$
Lemma 4.2.
Let $E\subset X$ be a compact set such that for some $r_{0},C>0$ we have
(4.1) $\frac{1}{C}r^{n}\leq\mathcal{H}^{n}(E\cap B(x,r))\leq Cr^{n}\qquad\forall r<r_{0},\,x\in E.$
Then for any $r<r_{0}$, there exists a cover $\mathcal{B}$ of $E$ by balls such that
  1. T1.
    the radius $r$ of every ball is the same;
  2. T2.
    $\sum_{B\in\mathcal{B}}\textnormal{diam}(B)\leq 8^{n}C\mathcal{H}^{n}(E)$;
  3. T3.
    for any $B\in\mathcal{B}$, #$\#B_{\cap}\leq C^{2}12^{n}$.
Proof.
Fix $r<r_{0}$ and let $\mathcal{C}=\{B(x_{i},r/2):x_{i}\in E\}_{i=1}^{N}$ be a finite covering of $E$, the existence of which is assured by the assumption of compactness. We first modify this covering such that the $x_{i}$ are distance at least $r/2$ apart. Iteratively pick balls such that the centre of one is not included in another and collect into a new collection $\mathcal{C}^{\prime}$ until no this is no longer possible. This terminates after finitely many steps. Set $\mathcal{B}=2\mathcal{C}^{\prime}$. We have that $\mathcal{B}$ covers $\mathcal{C}$ and is made of balls of radius $r$ with $r/2$-separated centres.
Now pick a ball $B\in\mathcal{B}$ and note that $\bigcup_{A\in B_{\cap}}A\subset 3B$ and that the balls formed taking the same centres but a quarter the radius, $B_{\cap}/4$, are disjoint. Thus, using (4.1),
$\frac{1}{C}(r/4)^{n}(\#B_{\cap})$ $\leq\mathcal{H}^{n}(E\cap\bigcup_{A\in B_{\cap}}A/4)$
$\leq\mathcal{H}^{n}(E\cap 3B)\leq C(3r)^{n}$
which gives T3..
To prove T2., note that, since the elements of $\mathcal{B}/4$ are disjoint,
$\sum_{B_{i}\in\mathcal{B}}\mathcal{H}^{n}(E\cap B(x_{i},r/4))=\mathcal{H}^{n}(E\cap\bigcup_{B\in\mathcal{B}}B/4)\leq\mathcal{H}^{n}(E).$
Also, by (4.1),
$(2r)^{n}=8^{n}C\left(\frac{1}{C}(r/4)^{n}\right)\leq 8^{n}C\mathcal{H}^{n}(E\cap B(x_{i},r/4)).$
Combining these two inequalities gives T2.. ∎
We now consider continua in metric space. Recall that a continuum $L\subset X$ is a compact connected set with $\mathcal{H}^{1}(L)<\infty$.
Lemma 4.3.
Let $F\subset X$ be a compact set, $l>0$ and $L^{i}$ be a sequence of continua in $X$ with $\mathcal{H}^{1}(L^{i})<l$ such that $\textnormal{dist}(F,x)\leq 1/i$ for all $x\in L^{i}$. Then there exists a subsequence and a continuum $L$ in $F$ such that
  1. (1)
    $\mathcal{H}^{1}(L)\leq l$;
  2. (2)
    for all $m\in\mathbb{N}$, there exists an $I>0$ such that (up to subsequence)
    $L_{\boldsymbol{\underline{\nicefrac{{1}}{{m}}}}}\supset L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i}\qquad\forall\,i>I.$
Proof.
By [1, Theorems 4.2.1 & 4.4.8], the $L^{i}$ may be covered by 1-Lipschitz images of $[0,2l]$. By [1, Theorem 4.4.3], there exists a subsequence of these maps converging uniformly to some $\gamma\colon[0,2l]\to F$. Setting $L=\gamma([0,l])$ gives the desired result. ∎
The following lemma is a variant of [3, Lemma 8]. In place of an ambient Euclidean space, we will use the Kuratowski embedding to isometrically embed a separable metric space $X$ into $\ell_{\infty}$, the space of real bounded sequences with the supremum norm. In order to use compactness arguments, we note that, if $F\subset\ell_{\infty}$ is compact, then so is $CC(F)$, the closed convex hull of $F$.
Lemma 4.4.
Let $X$ be a metric space and let $F\subset X$ be compact, $\mathcal{H}^{1}$-measurable and purely 1-unrectifiable with $\mathcal{H}^{1}(F)<\infty$. For $\varepsilon,\,l>0$, there exists a radius $r>0$ such that for any continuum $L$ in $X$ with $\mathcal{H}^{1}(L)<l$,
$\mathcal{H}^{1}(F\cap L_{\boldsymbol{\underline{r}}})<\varepsilon.$
Proof.
Identify $X$ with its isometric image in $\ell_{\infty}$ and suppose that the hypothesis does not hold for some $\varepsilon>0$. Then for all $i\in\mathbb{N}$, there exists a sequence of continua $L^{i}\subset X$ such that
$\mathcal{H}^{1}(L^{i})<l$
but
$\mathcal{H}^{1}(F\cap L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i})\geq\varepsilon.$
We may additionally assume that $\textnormal{dist}(CC(F),x)\leq 1/i$ for all $x\in L^{i}$. Indeed, since each $L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i}$ is convex, we may replace any part of $L^{i}$ that lies outside by the line segment connecting the end points of this part, creating a shorter curve. Thus we may apply Lemma 4.3 to find that there exists a continuum $L\subset CC(F)$, such that $\mathcal{H}^{1}(L)\leq l$ and such that (for a relabelled subsequence of $L^{i}$ and) all $m\in\mathbb{N}$, there exists an $I>0$ such that
$L_{\boldsymbol{\underline{\nicefrac{{1}}{{m}}}}}\supset L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i}\qquad\forall\,i>I.$
Thus we have
$\mathcal{H}^{1}(F\cap L_{\boldsymbol{\underline{\nicefrac{{1}}{{m}}}}})\geq\mathcal{H}^{1}(F\cap L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i})\geq\varepsilon.$
Then observing the fact that $L=\bigcap_{m}L_{\boldsymbol{\underline{\nicefrac{{1}}{{m}}}}}$, and that $\mathcal{H}^{1}(L)\leq l$, we may use the continuity of $\mathcal{H}^{1}$ from above to see that
(4.2) $\mathcal{H}^{1}(F\cap L)=\mathcal{H}^{1}(F\cap\bigcap_{m}L_{\boldsymbol{\underline{\nicefrac{{1}}{{m}}}}})=\inf\mathcal{H}^{1}(F\cap L_{\boldsymbol{\underline{\nicefrac{{1}}{{m}}}}})\geq\varepsilon.$
However, $L$ is 1-rectifiable, thereby (4.2) contradicts the fact that $F$ is purely $1$-unrectifiable. ∎
We replace the notion of Besicovitch circle pairs with the following stronger analogue:
Definition 4.5 (Besicovitch partition).
A Besicovitch partition of a set $E\subset X$ is a collection $P=(E_{1},E_{2},\pi,p_{1},p_{2})$ containing
  1. (i)
    A partition: $E=E_{1}\cup E_{2}$;
  2. (ii)
    A distance: $0<\pi=\textnormal{dist}(E_{1},E_{2})$;
  3. (iii)
    A pair of points: $p_{1}\in E_{1}$ and $p_{2}\in E_{2}$ such that $d(p_{1},p_{2})=\pi$.
We require a notion of disjointness of Besicovitch partitions at a scale corresponding to the $\delta$-path radius that we found via Lemma 3.7. In this vein, we say that two Besicovitch partitions $P=(E_{1},E_{2},\pi,p_{1},p_{2})$ and $P^{\prime}=(E^{\prime}_{1},E^{\prime}_{2},\pi^{\prime},p^{\prime}_{1},p^{\prime}_{2})$ are $R$-disjoint if $B(p_{i},R\pi)\cap B(p^{\prime}_{j},R\pi^{\prime})=\emptyset$ for $i,j\in\{1,2\}$.
The following result is in the spirit of [3, Lemma 9]. However, the construction of the continua in the proof differs quite significantly from the original, which uses the fact that, in the plane, the boundary of a ball is a continuum.
Proposition 4.6.
Let $(X,d)$ be a complete metric space and $F\subset X$ compact, purely $1$-unrectifiable such that for some $r_{0}>0$ and $C\geq 1$, we have
(4.3) $\frac{1}{C}r\leq\mathcal{H}^{1}(F\cap B(x,r))\leq Cr\qquad\forall r<r_{0},\,x\in F.$
For any $\varepsilon,\delta>0$ and $R\geq 1$, there exists a finite collection $\mathcal{P}$ of Besicovitch partitions $P_{i}=(F_{i,1}F_{i,2},\pi_{i},p_{i,1},p_{i,2})$, $1\leq i\leq n$, such that
  1. P1.
    $P_{i}$, $P_{j}$ are $R$-disjoint for each $1\leq i\neq j\leq n$;
  2. P2.
    for each $1\leq i\leq n$, $\pi_{i}<\delta$;
  3. P3.
    the distances are cumulatively large:
    $\sum_{1\leq i\leq n}\pi_{i}>\frac{1-\varepsilon}{3(2R+1)}\mathcal{H}^{1}(F).$
Proof.
We identify $F$ with its isometric image in $\ell_{\infty}$. Let $\delta_{0}>0$ be given by Lemma 2.2, such that, for all $r\in(0,\,\delta_{0})$,
(4.4) $\mathcal{H}^{1}(F\cap\bigcup\mathcal{A}_{r})<\sum_{A_{i}\in\mathcal{A}_{r}}\textnormal{diam}(A_{i})+\frac{\varepsilon}{2}\mathcal{H}^{1}(F),$
where $\mathcal{A}_{r}$ is any collection of convex sets with diameter at most $r$.
Let
$\delta_{1}<\frac{1}{3(2R+1)}\min\{\delta,\delta_{0},r_{0}\}.$
Since $F$ is compact, there exists a finite set $\{x_{n}\}_{n=1}^{N}\subset F$, such that $F\subset\bigcup_{n=1}^{N}B(x_{n},\delta_{1})$. Let $L$ be the polygonal line joining these $x_{n}$, then $\mathcal{H}^{1}(L)=\sum d(x_{n},x_{n+1})$. Let $r<\delta_{1}$ be given by Lemma 4.4 such that, for any continuum $\tilde{L}$ with
$\mathcal{H}^{1}(\tilde{L})<\mathcal{H}^{1}(L)+200C^{3}(R+2)\mathcal{H}^{1}(F),$
we have
(4.5) $\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
Finally, we satisfy the assumptions of Lemma 4.2 so way may find a cover of $F$ by collection of balls $\mathcal{A}_{1}$ that satisfies
  1. T1.
    the radius $\hat{r}<r/2$ of every ball is the same;
  2. T2.
    $\sum_{B\in\mathcal{A}_{1}}\textnormal{diam}(B)\leq 8C\mathcal{H}^{1}(F)$;
  3. T3.
    for any $B\in\mathcal{A}_{1}$, #$\#B_{\cap}\leq 12C^{2}$.
We now iteratively construct collections $\mathcal{A}_{k}$ of balls by adding elements to $\mathcal{A}_{1}$ until the union of some $\mathcal{A}_{k}$ is connected. Suppose that $\mathcal{A}$ is a collection of balls. For each $B\in\mathcal{A}$ define
$\operatorname{Cluster}(B)=\left\{\hat{B}\in\mathcal{A}:B,\hat{B}\text{ in sameconnected component of }\left(\bigcup\mathcal{A}\right)_{\boldsymbol{\underline{\hat{r}/4}}}\right\},$
and
$\text{Cluster}{\mathcal{A}}:=\left\{\operatorname{Cluster}(B):B\in\mathcal{A}\right\}.$
Also define
$\text{Base}{\mathcal{A}}:=\bigcup\{C\in\text{Cluster}\mathcal{A}:x_{n}\in C\text{ for some }n\}.$
We now add balls to $\mathcal{A}_{k}$ as follows. Set $\mathcal{P}_{0}=\mathcal{B}_{0}=\emptyset$ and, for $k\geq 1$, suppose that $\mathcal{A}_{k}$ is defined.
  1. (1)
    Let $F_{1}=F\cap\text{Base}{\mathcal{A}_{k}}$ and $F_{2}=F\backslash F_{1}$.
  2. (2)
    Find points $p_{1}\in F_{1}$ and $p_{2}\in F_{2}$ such that $\pi:=d(p_{1},p_{2})=\textnormal{dist}(F_{1},F_{2})$, so that $(F_{1},F_{2},\pi,p_{1},p_{2})$ is a Besicovitch partition that satisfies P2..
  3. (3)
    Let $b$ be the midpoint of $p_{1}$ and $p_{2}$ and let $B=B(b,\frac{2R+1}{2}\pi)$.
  4. (4)
    Set $\mathcal{P}_{k}=\mathcal{P}_{k-1}\cup\{(F_{1},F_{2},\pi,p_{1},p_{2})\}$, $\mathcal{B}_{k}=\mathcal{B}_{k-1}\cup\{B\}$ and $\mathcal{A}_{k+1}=\mathcal{A}_{k}\cup B$.
  5. (5)
    If $\text{Base}{\mathcal{A}_{k+1}}=\bigcup\mathcal{A}_{k+1}$ , stop. Otherwise, repeat with $k=k+1$.
$\delta_{1}$
$x_{n-1}$
$\text{Base}{\mathcal{A}_{k}}$
$x_{n}$
$x_{n}$
$x_{n+1}$
$p_{1}$
$p_{2}$
$b$
$\pi$
Figure 2. Step 2 of the proof of Proposition 4.6: Finding the ball in $\mathcal{B}_{k}$ that corresponds to $\text{Base}{\mathcal{A}_{k}}$.
Since the $A_{i}$ are convex, this process will terminate after $K\leq m$ steps. Finally, define $\mathcal{A}$ by removing from $\mathcal{A}_{K}$ those $A_{i}$ with $A_{i}\subset B$ for some $B\in\mathcal{B}_{K}$. Then $\mathcal{A}\cup\mathcal{B}_{K}$ forms a covering of $F$ such that each connected component of $\bigcup(\mathcal{A}\cup\mathcal{B}_{K})$ contains an $x_{i}$.
The collections $\mathcal{B}_{K}$ and $\mathcal{P}_{K}$ will now be refined slightly to obtain P1.. Reindex $\mathcal{B}_{K}$ (and hence $\mathcal{P}_{K}$) such that $\textnormal{diam}B_{i}\geq\textnormal{diam}B_{j}$ for all $1\leq i\leq j\leq K$ and define
:=$\mathcal{B}:=\{B_{i}\in\mathcal{B}_{K}:B_{i}\cap B_{j}=\emptyset\ \forall 1\leq j<i\}.$
Also define
$\mathcal{P}:=\{(F_{1},F_{2},\pi,p_{1},p_{2})_{k}\in\mathcal{P}_{K}:B_{k}\in\mathcal{B}\}.$
Then the partitions in $\mathcal{P}$ now satisfy P1., as required. We also have that $\mathcal{B}_{K}$ is covered by $3\mathcal{B}$ and thus $\mathcal{A}\cup 3\mathcal{B}$ forms a cover of $F$.
To complete the proof, P3. is implied by
(4.6) $\sum_{B_{i}\in\mathcal{B}}3\textnormal{diam}(B_{i})>(1-\varepsilon)\mathcal{H}^{1}(F).$
Suppose that this inequality is not true. To derive a contradiction, we first must build a particular continuum which we do in three steps.
  1. (1)
    For any $B\in\mathcal{A}$, connect the centre of $B$ to the centres of each ball in $B_{\cap}$. By T3. there are at most $12C^{2}$ such balls. Repeat this for every ball in $\mathcal{A}$ to obtain a union of continua, $L_{1}$. By T2., $L_{1}$ satisfies
    $\mathcal{H}^{1}(L_{1})\leq 12C^{2}\sum_{B\in\mathcal{A}}\textnormal{diam}(B)\leq 96C^{3}\mathcal{H}^{1}(F).$
    Note that, for each connected component of $\cup\mathcal{A}$, $L_{1}$ connects the centres of the balls forming that component.
  2. (2)
    For each $B\in\mathcal{A}$, consider the collection
    $\text{Nbours}(B):=\left\{\hat{B}\in\mathcal{A}:0<\textnormal{dist}(B,\hat{B})\leq\hat{r}/4\right\}.$
    These balls are disjoint and are covered by $4B$. Thus, using (4.3), we have that
    $\frac{1}{C}\hat{r}\cdot\#\text{Nbours}(B)\leq\sum_{\hat{B}\in\text{Nbours}(B)}\mathcal{H}^{1}(F\cap\hat{B})\leq\mathcal{H}^{1}(F\cap 4B)\leq 4C\hat{r}$
    and hence
    #$\#\text{Nbours}(B)\leq 4C^{2}.$
    Thus we may connect the centre of $B$ to the centre of each ball in $\text{Nbours}(B)$ using a continuum of length at most $4C^{2}\cdot 3\hat{r}$. Repeat for each $B\in\mathcal{A}$ to create a union $L_{2}$ of disjoint continua satisfying
    $\mathcal{H}^{1}(L_{2})\leq 6C^{2}\sum_{B\in\mathcal{A}}\textnormal{diam}(B)\leq 72C^{3}\mathcal{H}^{1}(F).$
    Note that, for each connected component of $(\cup\mathcal{A})_{\boldsymbol{\underline{\hat{r}/4}}}$, $L_{1}\cup L_{2}$ connects the centres of the balls forming that component.
  3. (3)
    For $B\in 3\mathcal{B}$, consider the elements $C_{1},\ldots,C_{N}$ of $\text{Cluster}\mathcal{A}$ that intersect $B$. For each $C_{n}$, pick a ball $B(a_{n},\hat{r})\in\mathcal{A}$ that is a subset of $A_{n}$ and that intersects $B$. Note that by construction of $B$ we necessarily have
    $d(a_{n},a_{k})\geq\pi,\qquad\forall n\neq k$
    and
    $B(a_{n},\pi/2)\subset B_{\boldsymbol{\underline{\hat{r}+\pi/2}}}.$
    Hence
    $\frac{N}{C}\pi/2\leq\sum_{n=1}^{N}\mathcal{H}^{1}(F\cap B(a_{n},\pi/2))$ $\leq\mathcal{H}^{1}(F\cap B_{\boldsymbol{\underline{\hat{r}+\pi/2}}})$
    $\leq C(3(2R+1)\pi/2+\hat{r}+\pi/2).$
    Since, by contruction, $\pi\geq\hat{r}/4$, this implies
    $N\leq C^{2}(3(2R+1)+2\hat{r}/\pi+1)\leq C^{2}(3(2R+1)+9).$
    Connect the centre of $B$ to each $a_{n}$ to create a continuum of length at most
    $C^{2}(3(2R+1)+9)\cdot(\pi/2+\hat{r})\leq 6C^{2}(3(2R+1)+9)\pi.$
    Repeat this for every $B\in 3\mathcal{B}$ to form a union of continua $L_{3}$ which satisfies
    $\mathcal{H}^{1}(L_{3})$ $\leq C^{2}(3(2R+1)+9)\sum_{B\in 3\mathcal{B}}3\textnormal{diam}(B)$
    $\leq 6C^{2}(R+2)(1-\varepsilon)\mathcal{H}^{1}(F).$
All together we have a continuum $\tilde{L}=L\cup L_{1}\cup L_{2}\cup L_{3}$ such that
$\mathcal{H}^{1}(\tilde{L})$ $\leq\mathcal{H}^{1}(L)+96C^{3}\mathcal{H}^{1}(F)+72C^{3}\mathcal{H}^{1}(F)+6C^{2}(R+2)(1-\varepsilon)\mathcal{H}^{1}(F)$
$<\mathcal{H}^{1}(L)+200C^{3}(R+2)\mathcal{H}^{1}(F).$
Thus, (4.5) gives
$\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
Since each element of $\mathcal{A}$ has diameter $\hat{r}<r$,
$\mathcal{H}^{1}(F\cap\bigcup\mathcal{A})\leq\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
Therefore, since $\mathcal{A}\cup 3\mathcal{B}$ covers $F$,
$\mathcal{H}^{1}(F\setminus\bigcup 3\mathcal{B})\leq\mathcal{H}^{1}(F\cap\bigcup\mathcal{A})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F)$
and so
(4.7) $\mathcal{H}^{1}(F\cap\bigcup 3\mathcal{B})=\mathcal{H}^{1}(F)-\mathcal{H}^{1}(F\backslash\bigcup 3\mathcal{B})>(1-\varepsilon/2)\mathcal{H}^{1}(F).$
However, the diameters of the balls in $3\mathcal{B}$ are less than $\delta_{0}$. Therefore, combining (4.4) and the opposite of (4.6) gives
(4.8) $\mathcal{H}^{1}(F\cap\bigcup 3\mathcal{B})<\sum_{B_{i}\in\mathcal{B}}3\textnormal{diam}(B_{i})+\frac{\varepsilon}{2}\mathcal{H}^{1}(F)<(1-\varepsilon/2)\mathcal{H}^{1}(F),$
which contradicts (4.7). ∎
5. Connected Tangents Implies Rectifiable
We now prove the main theorem. The proof is rather technical however the underlying geometric idea is simple: a set cannot both be approximated by quasi-path connected spaces (which is implied by having connected tangents) and have collections of Besicovitch partitions as in Proposition 4.6 (which is implied by being purely unrectifiable).
Theorem 5.1.
Let $(X,d)$ be a complete metric space and $E\subset X$ be $\mathcal{H}^{1}$-measurable with $\mathcal{H}^{1}(E)<\infty$. If for $\mathcal{H}^{1}$-a.e. $x\in E$, $\Theta_{*}(E,x)>0$ and
$\textnormal{Tan}(X,d,\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{C}^{*},$
then $E$ is $1$-rectifiable.
Proof.
We proceed by contradiction and assume $E$ is not $1$-rectifiable. First, by Lemma 2.5, we may find a purely $1$-unrectifiable set $E^{\prime}\subset E$ of positive measure. By Corollary 2.8, we have $\Theta_{*}(E^{\prime},x)>0$ for $\mathcal{H}^{1}$-a.e. $x\in E^{\prime}$, thus $\mathcal{H}^{1}\llcorner E^{\prime}$ is asymptotically doubling and by Lemma 2.29,
$\textnormal{Tan}(\mathcal{H}^{1}\llcorner E^{\prime},x)=\textnormal{Tan}(\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{C}^{*}\quad\text{for }\mathcal{H}^{1}\text{-a.e.}\,x\in E^{\prime}.$
Therefore we satisfy the requirements of Lemma 3.10. Thus, for $\delta=1/6$, we may find a set $F\subset E^{\prime}$ and numbers $\hat{d},R>0$ such that for $\mathcal{H}^{1}$-a.e. $x\in F$,
S1. $\quad\hat{d}\leq\Theta_{*}(F,x),\,\Theta^{*}(F,x)\leq 1;$
S2. $\quad\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,x)\subset\mathcal{Q}_{*}(1/6,R,1/\hat{d})\cap\mathcal{C}^{*}.$
By the virtue of the fact that $Q_{*}$ are nested, $\mathcal{Q}_{*}(\delta,R,C)\subset\mathcal{Q}_{*}(\delta,R+1,C)$, we can assume that $R$ is such that $128/(R+1)<\hat{d}/$ 12. We may then find a further subset $G\subset F$ with
(5.1) $\mathcal{H}^{1}(F\backslash G)<\frac{\hat{d}^{2}(1-\varepsilon)}{72(2R+3)+1}\mathcal{H}^{1}(F)$
and an $r_{0}>0$ such that for all $x\in G$ and $r<r_{0}$
(5.2) $\hat{d}r\leq\mathcal{H}^{1}(F\cap B(x,r))\leq 4r,$
and such that, by Proposition 2.30,
(5.3) $d_{*}(T_{r}(\mathcal{H}^{1}\llcorner F,d,x),\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,x))<\frac{1}{R+1}.$
Fix $\varepsilon>0$ and $\delta\leq\min\{r_{0},\hat{d}\}/2$. Use Lemma 4.6 to find a collection $\mathcal{P}$ of Besicovitch partitions for $G$ satisfying
[…]
P3.$\quad\sum_{i}\pi_{i}>\frac{1-\varepsilon}{3(2R+3)}\mathcal{H}^{1}(G).$
For a moment fix a $P=(G_{1},G_{2},\hat{\pi},p_{1},p_{2})$ in $\mathcal{{P}}$ and set $\pi=2\hat{\pi}$. Because $\pi<r_{1}$, we can find $(\nu,\rho,y)\in\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,p_{1})$ such that (5.3) holds for $r=\pi$. Then by Theorem 2.26, there exists compact $p_{1}\in K_{F}\subset F\cap B(p_{1},\pi(R+1))$ and $y\in K_{\nu}\subset\operatorname{supp}\nu\cap B(y,R+1)$ such that
M1. $\max\left\{\frac{\mathcal{H}^{1}(F\cap B(p_{1},\pi(R+1))\backslash K_{F})}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},\nu(B(y,R+1)\backslash K_{\nu})\right\}<\frac{1}{R+1},$
M2. $d_{pmGH}\left((K_{F},\frac{\mathcal{H}^{1}\llcorner K_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},p_{1}),(K_{\nu},\nu\llcorner K_{\nu},y)\right)<\frac{3}{R+1}.$
By M2. and Proposition 2.27, there exists a metric space $(Z,\lambda)$ and isometric embeddings
$(K_{F},p_{1}),(K_{\nu},y)\rightarrow(\bar{K}_{F},z),(\bar{K}_{\nu},z)\subset(Z,\lambda,z)$
such that
(5.4) $F_{z}\left(\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},\bar{\nu}\llcorner\bar{K}_{\nu}\right)+d_{H(z)}(\bar{K}_{F},\bar{K}_{\nu})<\frac{3}{R+1},$
where for clarity and ease of notation, the image of any set or point under these embeddings is denoted with a bar.
We now demonstrate the existence of points in both $\bar{K}_{F}$ and $\bar{K}_{\mu}$ near $\bar{p}_{2}$.
Firstly, note that by our choice of $R$ we now satisfy the assumptions of Proposition 3.4 where $\varepsilon=3/(R+1)$ and $\delta=1/6$, thus there exists $s\in K_{F}\cap B_{d/\pi}(p_{2},1/6)$. This point combined with (5.4), gives us the existence of $\bar{t}\in\bar{K}_{\nu}$ such that $\lambda(\bar{s},\bar{t})\leq 3/(R+1)$. Thus, using the triangle inequality we find that
$\lambda(z,\bar{t})$ $\leq\lambda(z,\bar{p}_{2})+\lambda(\bar{p}_{2},\bar{s})+\lambda(\bar{s},\bar{t})$
$=d_{d/\pi}(p_{1},p_{2})+d_{d/\pi}(p_{2},s)+\lambda(s,t)$
$\leq 1/2+1/6+3/(R+1)\leq 1.$
Thus we have that $t\in K_{\nu}$ is such that $\rho(y,t)\leq 1$. Once again, due to our choice of $R$, we satisfy the assumptions of Corollary 3.5, with $\varepsilon=3/(R+1)$ and $\delta=1/6$, thus we have that $K_{\nu}\cap B(y,R+1/12)$ contains a $1/3$-quasi-path connecting $y$ and $t$. Let $K_{G_{1}}=K_{F}\cap G_{1}$ and $K_{G_{2}}=K_{F}\cap G_{2}$, then we have $\textnormal{dist}(\bar{K}_{G_{1}},\,\bar{K}_{G_{2}})\geq 1/2.$ Because $\bar{K}_{\nu}$ contains a $1/3$-quasi-path between $z$ and $\bar{t}$, there must be some $\bar{q}\in\bar{K}_{\nu}$ such that
$\textnormal{dist}(\bar{K}_{G_{1}},\bar{q})\geq\frac{1}{6}\quad\text{and}\quad\textnormal{dist}(\bar{K}_{G_{2}},\bar{q})\geq\frac{1}{6}$
and $B(\bar{q},1/12)\subset B(z,R+1/24)\subset B(z,R+1)$. Thus there exists the required points near $\bar{p}_{2}$.
Now, since $\nu\in\mathcal{A}(1/\hat{d})$,
(5.5) $\nu(B(q,1/12)\cap K_{\nu})\geq\frac{\hat{d}}{12}-\frac{1}{R+1}$
which combined with (5.4) implies that
(5.6) $\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}(B(\bar{q},1/12))}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))}\geq\nu(B(\bar{q},1/12)\cap K_{\nu})-\frac{3}{R+1}\geq\frac{\hat{d}}{12}-\frac{4}{R+1}.$
Therefore,
$\mathcal{H}^{1}(F\cap B_{d/\pi}(p_{1},R/2)\backslash G_{1}\cup G_{2})$ $\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\mathcal{H}^{1}(F\cap B(p_{1},\pi))$
$\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\hat{d}\pi>\frac{\hat{d}^{2}}{24}\pi.$
Since the Besicovitch partitions are disjoint (condition P1.), the balls $B_{d/\pi}(p_{1},R/2)$ corresponding to each partition are disjoint. we may repeat the same process for each partition $P_{i}\in\mathcal{P}$ and use P3. to obtain
$\mathcal{H}^{1}(F\backslash G)$ $\geq\frac{\hat{d}^{2}}{24}\cdot\frac{1-\varepsilon}{3(2R+3)}\mathcal{H}^{1}(G)$
$=\frac{\hat{d^{2}}(1-\varepsilon)}{72(2R+3)}(\mathcal{H}^{1}(F)-\mathcal{H}^{1}(F\backslash G)).$
This implies
$\mathcal{H}^{1}(F\backslash G)\geq\frac{\hat{d}^{2}(1-\varepsilon)}{72(2R+3)+\hat{d}^{2}(1-\varepsilon)}\mathcal{H}^{1}(F)\geq\frac{\hat{d}^{2}(1-\varepsilon)}{72(2R+3)+1}\mathcal{H}^{1}(F)$
which contradicts (5.1) and concludes the proof. ∎
Finally, we restate the necessary part of [2, Theorem 1.2], with which we obtain our final statement. For $K\geq 1$, let $\textnormal{biLip}(K)^{*}$ be the set of all pointed metric measure spaces whose metric is $K$-bi-Lipschitz equivalent to the Euclidean norm.
Theorem 5.2 (Theorem 1.2 [2]).
Let $(X,d)$ be a complete metric space, $n\in\mathbb{N}$ and let $E\subset X$ be $\mathcal{H}^{n}$-measurable with $\mathcal{H}^{n}(E)<\infty$. Then $E$ is $n$-rectifiable if and only if for $\mathcal{H}^{n}$-a.e. $x\in E$, $\Theta_{*}^{n}(E,x)>0$ and there exists a $K_{x}\geq 1$ such that
$\textnormal{Tan}(X,d,\mathcal{H}^{n}\llcorner E,x)\subset\textnormal{biLip}(K_{x})^{*}.$
Proof of Theorem 1.1.
We have that $\text{biLip}(K)^{*}$ is a subset of $\mathcal{C}^{*}$ so simply combine Theorem 5.2 and Lemma 5.1. ∎

### v2

@@
 Abstract.
-We prove that in a complete metric space $X$, $1$-rectifiability of a set $E\subset X$ with $\mathcal{H}^{1}(E)<\infty$ and positive lower density $\mathcal{H}^{1}$-a.e.
+We prove that in a complete metric space $X$, 11-rectifiability of a set $E\subset X$ with $\mathcal{H}^{1}(E)<\infty$ and positive lower density $\mathcal{H}^{1}$-a.e.
 is implied by the property that all tangent spaces are connected metric spaces.
@@
 For a metric space $(X,d)$, a set $E\subset X$ is $n$-rectifiable if there exists countably many Lipschitz maps $f_{i}\colon A_{i}\subset\mathbb{R}^{n}\rightarrow X$ such that
-$\mathcal{H}^{n}(E\backslash\bigcup_{i\in\mathbb{N}}f_{i}(A_{i}))=0.$
+[…]
 Here and throughout, $\mathcal{H}^{n}$ denotes the $n$-dimensional Hausdorff (outer) measure.
@@
 In this article we show that, for 1-dimensional sets, this sufficient condition for rectifiability can be weakened further to tangents that are supported on connected metric spaces.
-Let $\mathcal{C}_{*}$ be the space of pointed metric measure spaces $(X,d,\mu,x)$ for which the support of $\mu$ is connected.
+Let $\mathcal{C}^{*}$ be the space of pointed metric measure spaces $(X,d,\mu,x)$ for which the support of $\mu$ is connected.
 Theorem 1.1.
@@
     $\textnormal{Tan}(X,d,\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{C}^{*}.$
-The fact that 1) implies 2) is given in [2, Theorem 6.6] by applying the results of Kirchheim [5].
+The fact that 1) implies 2) is given in [3, Theorem 6.6] by applying the results of Kirchheim [6].
 Since 2) implies 3) is immediate, our attention is focused on the implication 3) implies 1).
-We emphasise that, even in the case that all tangent spaces are supported on $\mathbb{R}$, our proof follows a completely different approach that of [2].
-Indeed, the geometric idea of our proof begins with a result o Besicovitch [3, Lemma 9] that covers a purely 1-unrectifiable subset of $\mathbb{R}^{2}$ by a large number of circle pairs.
+We emphasise that, even in the case that all tangent spaces are supported on $\mathbb{R}$, our proof follows a completely different approach than that of [3].
+Indeed, the geometric idea of our proof begins with a result of Besicovitch [4, Lemma 9] that covers a purely 1-unrectifiable subset of $\mathbb{R}^{2}$ by a large number of circle pairs.
 A circle pair of a set $F$ is simply $B(x,|x-y|)\cup B(y,|x-y|)$, where $x$ and $y$ are points in $F$.
@@
 These results, and the progress towards the conjecture, were recently extended further by De Lellis, Glaudo, Massaccesi and Vittone [7].
+Further, our result can be viewed as a substantial strengthening of a theorem of Azzam and Mourgoglou [2], which states that a doubling metric measure space with positive 1-density and connected support is 1-rectifiable.
 The structure of the paper is as follows.
@@
 $\mathcal{H}^{s}(E):=\lim_{\delta\rightarrow 0}\mathcal{H}^{s}_{\delta}(E).$
-For any $s$, Hausdorff measure is a Borel regular measure [10, Corollary 4.5].
+For any $s$, Hausdorff measure is a Borel regular measure [11, Corollary 4.5].
 Lemma 2.2.
@@
 Let $\{A_{i}\subset X:\textnormal{diam}(A_{i})\leq\delta_{0}\}_{i\in\mathbb{N}}$, then
-$\mathcal{H}^{s}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})+\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})$ $=\mathcal{H}^{s}(E)$
+[…]
 $\leq\mathcal{H}^{s}_{\delta}(E)+\varepsilon$
-$=\mathcal{H}^{s}_{\delta}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})+\mathcal{H}^{s}_{\delta}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})+\varepsilon$
-$\leq\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})+\sum_{i}\textnormal{diam}(A_{i})^{s}+\varepsilon.$
-Subtracting $\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})$ from both sides (which is finite by assumption) gives the desired result. ∎
+[…]
+[…]
+[…]
 Definition 2.3 (Lipschitz Map).
-Given a metric space $(X,d)$ and a set $A\subset\mathbb{R}^{n}$, a map $f:A\rightarrow X$ is Lipschitz if there exists $K>0$ such that
+Given a metric space $(X,d)$ and a set $A\subset\mathbb{R}^{n}$, a map $f:A\rightarrow X$ is Lipschitz if there exists $K\geq 0$ such that
 $d(f(x),f(y))\leq K|x-y|,\qquad\forall x,\,y\in A.$
@@
 Proof.
-The proof for $X=\mathbb{R}^{m}$ is given in [10, Theorem 15.6]. The proof for general $X$ is identical. ∎
+The proof for $X=\mathbb{R}^{m}$ is given in [11, Theorem 15.6]. The proof for general $X$ is identical. ∎
 2.3. Density and Doubling
@@
   2. (ii)
-    If $E$ is $\mathcal{H}^{s}$-measurable, then $\Theta^{*s}(E,x)=0$ for $\mathcal{H}^{s}$-almost all $x\in X\backslash E.$
+    If $E$ is $\mathcal{H}^{s}$-measurable, then $\Theta^{*s}(E,x)=0$ for $\mathcal{H}^{s}$-almost all $\E.x\in X\backslash$ E.
 Corollary 2.8.
@@
 For $K\geq 1$, $\mu\in\mathcal{M}(X)$ is $K$-doubling if
-[…]
+$0<\mu(B(x,\,2r))\leq K\mu(B(x,\,r))<\infty,\qquad\forall\,x\in X,\,r>0,$
 or asymptotically doubling if
@@
 Definition 2.13 (Pointed metric (measure) space).
-A pointed metric space $(X,\,d,\,x)$ is an ordered triple consisting of a metric space $(X,\,d)$ and a point $x\in X$.
+A pointed metric space (X,d,x)(X,\,d,\,x) is an ordered triple consisting of a metric space (X,d)(X,\,d) and a point $x\in X$.
 A pointed metric measure space $(X,\,d,\,\mu,\,x)$ consists of a complete and separable pointed metric space $(X,x)$ and a measure $\mu\in\mathcal{M}_{loc}(X)$.
-An isometric embedding $(X,\,d,\,x)\rightarrow(Y,\,\rho,\,y)$ of pointed metric spaces is an isometric embedding $(X,\,d)\rightarrow(Y,\,\rho)$ such that $x=y.$
+An isometric embedding $(X,\,d,\,x)\rightarrow(Y,\,\rho,\,y)$ of pointed metric spaces is an isometric embedding $(X,\,d)\rightarrow(Y,\,\rho)$ such that $x\mapsto y.$
 Notation 2.14.
@@
 Let $(Z,d)$ be a metric space and $X,\,Y\subset Z$. Then for $z\in Z$, we define the pointed Hausdorff distance between $X$ and $Y$ as
-[…]
+(2.16) $d_{H(z)}(X,\,Y):=\inf\left\{\varepsilon>0:X\cap B(z,\,1/\varepsilon)\subset Y_{\underline{\varepsilon}},\,Y\cap B(z,\,1/\varepsilon)\subset X_{\underline{\varepsilon}}\right\}.$
 We can now define the pointed Gromov-Hausdorff distance which allows us to compare sets across different ambient metric spaces.
-Notation 2.16.
-In general we shal not distinguish between the original objects $X$ and $\mu$ and their isometric images $\iota(X)\subset Z$ and $\iota_{\#}\mu$.
+Notation 2.17.
+In general we shall not distinguish between the original objects $X$ and $\mu$ and their isometric images $\iota(X)\subset Z$ and $\iota_{\#}\mu$.
 Definition 2.18 (Pointed Gromov-Hausdorff distance).
-Let $(X,\,x)$ and $(Y,\,y)$ be pointed metric spaces. We define the pointed Gromov-Hausdorff distance as
+Let (X,x)(X,\,x) and (Y,y)(Y,\,y) be pointed metric spaces. We define the pointed Gromov-Hausdorff distance as
 […]
@@
 If $\mathbb{M}_{p}$ is the set of isometry classes of proper pointed metric spaces, then $(\mathbb{M}_{p},\,d_{pGH})$ is a complete and seperable metric space.
-In particular, if $(X,\,x)$ and $(Y,\,y)$ are proper with
+In particular, if (X,x)(X,\,x) and (Y,y)(Y,\,y) are proper with
 $d_{pGH}((X,x),(Y,y))=0$
@@
 A sequence $\mu_{i}\in\mathcal{M}_{loc}(X)$ weak* converges to $\mu\in\mathcal{M}_{loc}(X)$ if
+$\int f\,d\mu_{i}\rightarrow\int f\,d\mu$
 […]
-for all $f\in C_{bs}(X)$. We use $\xrightharpoonup{*}$ to denote this convergence.
 Proposition 2.21 ([8] Proposition 4.26).
@@
   3. (3)
-    If $B(x,r)$ is such that with $\mu(B(x,r)\backslash U(x,r))=0$, then
+[…]
     $\mu(B(x,r))=\lim_{i\rightarrow\infty}\mu_{i}(B(x_{i},r)).$
@@
 Definition 2.25.
-Let $x\in X$ and $\mu,\nu\in\mathcal{M}(X)$. For $L,\,r>0$ define
+Let $x\in X$ and $\mu,\nu\in\mathcal{M}(X)$. For L,r>0L,\,r>0 define
 $F^{L,\,r}_{x}(\mu,\nu):=\sup\left\{\int f\,d(\mu-\nu):f\in\textnormal{Lip}_{L}(X;[-1,1])\,\textnormal{ with }\operatorname{supp}f\subset B(x,\,r)\right\}.$
@@
 […]
-This is a complete and separable metric on the set of isometry classes of pointed metric measure spaces $\mathbb{M}_{pm}$, see [2, Corollary 4.6].
+This is a complete and separable metric on the set of isometry classes of pointed metric measure spaces $\mathbb{M}_{pm}$, see [3, Corollary 4.6].
 We are now ready to introduce the metric that will be used to define tangent measures in the next section.
@@
 $(X,\mu,x)\sim(Y,\nu,y)\text{ if }(\operatorname{supp}\mu\cup\{x\},\mu,x)\textnormal{ is isometric to }(\operatorname{supp}\nu\cup\{y\},\nu,y).$
-Then as shown in [2, Corollary 4.13], $(\mathbb{M}_{*},d_{*})$ is a complete and separable metric space.
+Then as shown in [3, Corollary 4.13], $(\mathbb{M}_{*},d_{*})$ is a complete and separable metric space.
 In light of this, to simplify notation we will sometimes write $(\mu,x)$ in place of $(X,\mu,x)$.
@@
   1. M1.
-    $\max\{\mu(B(x,1/\varepsilon)\backslash K_{\mu}),\nu(B(y,1/\varepsilon)\backslash K_{\nu})\}<\varepsilon$;
+    $max\{\mu(B(x,1/\varepsilon)\backslash K_{\mu}),\nu(B(y,1/\varepsilon)\backslash K_{\nu})\}<\varepsilon$;
   2. M2.
@@
 Tangent measures were first introduced by Preiss [12] in Euclidean space as a natural setting in which to explore questions around density.
-We refer the reader to [10, Page 184] for an introduction to tangent measures.
+We refer the reader to [11, Page 184] for an introduction to tangent measures.
 With the notion of convergence of metric measure spaces established in the previous section, the definition of tangent measures in metric spaces follows analogously.
@@
 The set of all tangent measures for $(\mu,d,x)$ is denoted $\textnormal{Tan}(\mu,d,x)$.
+Note that this definition of a tangent measure allows for different sequences of radii to produce different tangents - in a sense allowing the tangents to “rotate”.
+A basic example is that of a spiral in the plane that converges very quickly in the radial direction.
+In this case, the set of tangent measures at the origin consists of all half lines, equipped with Lebesgue measure.
 Corollary 2.32 (Corollary 4.7 [3]).
@@
 Let $X$ be a metric space and let $\mu\in\mathcal{M}_{loc}(X)$ be asymptotically doubling. For $\mu$-a.e. $x\in\operatorname{supp}\mu$ and every $r_{0}>0,$
-[…]
+$\{T_{r}(\mu,x):0<r<r_{0}\}\textnormal{ is pre-compact.}$
 In particular, for any $x\in\operatorname{supp}\mu$ satisfying the above, $\textnormal{Tan}(\mu,x)$ is a non-empty compact metric space when equipped with $d_{*}$ and
@@
     For $\mu$-a.e. $x\in\operatorname{supp}\mu$ and $s>0$, there exists $C>0$ such that
-    $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)\quad\implies\quad(C\nu,s\rho,y)\in\textnormal{Tan}(\mu,x)$
+    $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)\quad\implies\quad(C\nu,s\rho,y)\in\textnormal{Tan}(\mu,x).$
   2. (2)
@@
 Proof.
-Proof of (2) may be found in [2, Proposition 5.5] so we just present a proof of (1). If $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)$ then there exists $r_{i}\rightarrow 0$ such that
+Proof of (2) may be found in [3, Proposition 5.5] so we just present a proof of (1). If $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)$ then there exists $r_{i}\rightarrow 0$ such that
 $d_{*}(T_{r_{i}}(\mu,d,x),(\nu,\rho,y))\rightarrow 0.$
@@
 Suppose $(X,d,\mu,a)\notin\mathcal{Q}_{*}(\delta,\,R)$. Then there exists a point $b\in\operatorname{supp}\mu$ such that there is no $\delta$-quasi-path between $a$ and $b$ in $B(a,Rd(a,b))$. Define
-$A=\{x\in\operatorname{supp}\mu\cap B(a,Rd(a,b)):\\ \textnormal{there exists a $\delta$-quasi-path between $a$ and $x$ in $B(a,Rd(a,b))$}\}$
+[…]
 and $B=\operatorname{supp}\mu\cap B(a,Rd(a,b))\setminus A$. Then $a\in A$ and $b\in B$.
@@
 (3.9) $\mu(B(x_{n},\delta d(x,y)))>\varepsilon\qquad\forall\,n\in\mathbb{N}.$
-Then for any $K_{\mu}\subset X$ containing $x,\,y$ and satisfying
+Then for any $K_{\mu}\subset X$ containing x,yx,\,y and satisfying
 (3.10) $\mu(B(y,(R+1)d(x,y))\backslash K_{\mu})<\varepsilon,$
@@
 Proposition 3.12.
-Let $(X,\,d)$ be a complete metric space and $\mu\in\mathcal{M}_{loc}(X)$. Suppose that for $\mu$-a.e. $x\in\operatorname{supp}\mu$ we have
+Let (X,d)(X,\,d) be a complete metric space and $\mu\in\mathcal{M}_{loc}(X)$. Suppose that for $\mu$-a.e. $x\in\operatorname{supp}\mu$ we have
   1. (i)
@@
 Then by Proposition 2.26 we have $\mu_{i}\xrightharpoonup{*}\mu$.
-Let $y\in\operatorname{supp}\mu$. We will demonstrate that there exists a $\delta$-quasi-path between $x$ and $y$.
+Let $y\in\operatorname{supp}\mu$.
+We will demonstrate that there exists a $\delta$-quasi-path between $x$ and $y$.
 First, by Proposition 2.24, we know that $\mu\in\mathcal{A}(C)$.
@@
 $\mu(B)\geq\limsup_{i\rightarrow\infty}\mu_{i}(B)>0$
-which directly contradicts item (1), concluding the proof.
+which directly contradicts item (1), concluding the proof. ∎
 Corollary 3.16.
@@
 For any $R>0$, define the set
-$F_{R}=\{x\in E:\textnormal{Tan}(\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{Q}_{*}(\delta,R,1/\hat{d})\}.$
+[…]
 By Proposition 3.12 and Proposition 2.35, we know that $F_{R}$ monotonically increase with $R$ to cover $\mathcal{H}^{1}$-almost all of $E^{\prime}$.
@@
 $\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,x)=\textnormal{Tan}(\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{Q}_{*}(\delta,R,1/\hat{d})\cap\mathcal{C}^{*},$
-which gives us S2. as required.
+which gives us S2. as required. ∎
 4. Besicovitch Partitions
@@
 Indeed, in the next section we will see that a set with a suitable collection of Besicovitch partitions cannot have quasi-path connected tangents.
-In place of the Besicovitch covering theorem, we use the Alhfors type estimate (4.1) to bound the size of the cover.
+In place of the Besicovitch covering theorem, we use the Ahlfors type estimate (4.3) to bound the size of the cover.
 Notation 4.1.
-For a collection $\mathcal{A}$ and a set $A\in\mathcal{A}$, let
-$A_{\cap}:=\{B\in\mathcal{A}:A\cap B\neq\emptyset\}.$
+For a collection $\mathcal{B}$ and a set $B\in\mathcal{B}$, let
+$B_{\cap}:=\{B^{\prime}\in\mathcal{B}:B\cap B^{\prime}\neq\emptyset\}.$
 Lemma 4.2.
-Let $E\subset X$ be a compact set such that for some $r_{0},C>0$ we have
+Let $E\subset X$ be a totally bounded set such that for some $r_{0},C>0$ we have
 (4.3) $\frac{1}{C}r^{n}\leq\mathcal{H}^{n}(E\cap B(x,r))\leq Cr^{n}\qquad\forall r<r_{0},\,x\in E.$
-Then for any $r<r_{0}$, there exists a cover $\mathcal{B}$ of $E$ by balls such that
+Then for any $r<r_{0}$, there exists a finite cover $\mathcal{B}$ of $E$ by balls such that
   1. T1.
@@
   3. T3.
-    for any $B\in\mathcal{B}$, #$\#B_{\cap}\leq C^{2}12^{n}$.
+    for any $B\in\mathcal{B}$, $B_{\cap}\leq C^{2}12^{n}$.
 Proof.
@@
 Combining these two inequalities gives T2.. ∎
-We now consider continua in metric space. Recall that a continuum $L\subset X$ is a compact connected set with $\mathcal{H}^{1}(L)<\infty$.
+We now consider continua in metric space.
+Recall that a continuum $L\subset X$ is a compact connected set with $\mathcal{H}^{1}(L)<\infty$.
+Recall also that a set with an underlined subscript denotes a closed neighbourhood that set:
+$L_{\boldsymbol{\underline{\varepsilon}}}:=\{x\in X:\textnormal{dist}(x,L)\leq\varepsilon\}.$
 Lemma 4.4.
@@
 By [1, Theorem 4.4.3], there exists a subsequence of these maps converging uniformly to some $\gamma\colon[0,2l]\to F$.
-Setting $L=\gamma([0,l])$ gives the desired result.
+Setting $L=\gamma([0,2l])$ gives the desired result.
 ∎
-The following lemma is a variant of [3, Lemma 8].
+The following lemma is a variant of [4, Lemma 8].
 In place of an ambient Euclidean space, we will use the Kuratowski embedding to isometrically embed a separable metric space $X$ into $\ell_{\infty}$, the space of real bounded sequences with the supremum norm.
@@
 We may additionally assume that $\textnormal{dist}(CC(F),x)\leq 1/i$ for all $x\in L^{i}$.
-Indeed, since each $L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i}$ is convex, we may replace any part of $L^{i}$ that lies outside by the line segment connecting the end points of this part, creating a shorter curve.
+Indeed, because $CC(F)$ is convex, we may replace any part of $L^{i}$ that lies outside by the line segment connecting the end points of this part, creating a shorter curve.
 Thus we may apply Lemma 4.4 to find that there exists a continuum $L\subset CC(F)$, such that $\mathcal{H}^{1}(L)\leq l$ and such that (for a relabelled subsequence of $L^{i}$ and) all $m\in\mathbb{N}$, there exists an $I>0$ such that
@@
 Definition 4.7 (Besicovitch partition).
-A Besicovitch partition of a set $E\subset X$ is a collection $P=(E_{1},E_{2},\pi,p_{1},p_{2})$ containing
+A Besicovitch partition of a set $E\subset X$ is a collection $P=(E_{1},E_{2},\omega,p_{1},p_{2})$ containing
   1. (i)
@@
   2. (ii)
-    A distance: $0<\pi=\textnormal{dist}(E_{1},E_{2})$;
+    A distance: $0<\omega=\textnormal{dist}(E_{1},E_{2})$;
   3. (iii)
-    A pair of points: $p_{1}\in E_{1}$ and $p_{2}\in E_{2}$ such that $d(p_{1},p_{2})=\pi$.
+    A pair of points: $p_{1}\in E_{1}$ and $p_{2}\in E_{2}$ such that $d(p_{1},p_{2})=\omega$.
 We require a notion of disjointness of Besicovitch partitions at a scale corresponding to the $\delta$-path radius that we found via Lemma 3.12.
-In this vein, we say that two Besicovitch partitions $P=(E_{1},E_{2},\pi,p_{1},p_{2})$ and $P^{\prime}=(E^{\prime}_{1},E^{\prime}_{2},\pi^{\prime},p^{\prime}_{1},p^{\prime}_{2})$ are $R$-disjoint if $B(p_{i},R\pi)\cap B(p^{\prime}_{j},R\pi^{\prime})=\emptyset$ for $i,j\in\{1,2\}$.
-The following result is in the spirit of [3, Lemma 9].
+In this vein, we say that two Besicovitch partitions $P=(E_{1},E_{2},\omega,p_{1},p_{2})$ and $P^{\prime}=(E^{\prime}_{1},E^{\prime}_{2},\omega^{\prime},p^{\prime}_{1},p^{\prime}_{2})$ are $R$-disjoint if $B(p_{i},R\omega)\cap B(p^{\prime}_{j},R\omega^{\prime})=\emptyset$ for $i,j\in\{1,2\}$.
+The following result is in the spirit of [4, Lemma 9].
 However, the construction of the continua in the proof differs quite significantly from the original, which uses the fact that, in the plane, the boundary of a ball is a continuum.
@@
 (4.9) $\frac{1}{C}r\leq\mathcal{H}^{1}(F\cap B(x,r))\leq Cr\qquad\forall r<r_{0},\,x\in F.$
-For any $\varepsilon,\delta>0$ and $R\geq 1$, there exists a finite collection $\mathcal{P}$ of Besicovitch partitions $P_{i}=(F_{i,1}F_{i,2},\pi_{i},p_{i,1},p_{i,2})$, $1\leq i\leq n$, such that
+For any $\varepsilon,\delta>0$ and $R\geq 1$, there exists a finite collection $\mathcal{P}$ of Besicovitch partitions $P_{i}=(F_{i,1}F_{i,2},\omega_{i},p_{i,1},p_{i,2})$, $1\leq i\leq n$, such that
   1. P1.
@@
   2. P2.
-    for each $1\leq i\leq n$, $\pi_{i}<\delta$;
+    for each $1\leq i\leq n$, $\omega_{i}<\delta$;
   3. P3.
     the distances are cumulatively large:
-    $\sum_{1\leq i\leq n}\pi_{i}>\frac{1-\varepsilon}{3(2R+1)}\mathcal{H}^{1}(F).$
+    $\sum_{1\leq i\leq n}\omega_{i}>\frac{1-\varepsilon}{3(2R+1)}\mathcal{H}^{1}(F).$
 Proof.
@@
 we have
-(4.5) $\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
-Finally, we satisfy the assumptions of Lemma 4.2 so way may find a cover of $F$ by collection of balls $\mathcal{A}_{1}$ that satisfies
+[…]
+Finally, we satisfy the assumptions of Lemma 4.2 so way may find a cover of $F$ by a finite collection of balls $\mathcal{A}_{1}$ that satisfies
   1. T1.
@@
   3. T3.
-    for any $B\in\mathcal{A}_{1}$, #$\#B_{\cap}\leq 12C^{2}$.
+    for any $B\in\mathcal{A}_{1}$, $B_{\cap}\leq 12C^{2}$.
 We now iteratively construct collections $\mathcal{A}_{k}$ of balls by adding elements to $\mathcal{A}_{1}$ until the union of some $\mathcal{A}_{k}$ is connected.
@@
   2. (2)
-    Find points $p_{1}\in F_{1}$ and $p_{2}\in F_{2}$ such that $\pi:=d(p_{1},p_{2})=\textnormal{dist}(F_{1},F_{2})$, so that $(F_{1},F_{2},\pi,p_{1},p_{2})$ is a Besicovitch partition that satisfies P2..
+    Find points $p_{1}\in F_{1}$ and $p_{2}\in F_{2}$ such that $\omega:=d(p_{1},p_{2})=\textnormal{dist}(F_{1},F_{2})$, so that $(F_{1},F_{2},\omega,p_{1},p_{2})$ is a Besicovitch partition that satisfies P2..
   3. (3)
-    Let $b$ be the midpoint of $p_{1}$ and $p_{2}$ and let $B=B(b,\frac{2R+1}{2}\pi)$.
+    Let $b$ be the midpoint of $p_{1}$ and $p_{2}$ and let $B=B(b,\frac{2R+1}{2}\omega)$.
   4. (4)
-    Set $\mathcal{P}_{k}=\mathcal{P}_{k-1}\cup\{(F_{1},F_{2},\pi,p_{1},p_{2})\}$, $\mathcal{B}_{k}=\mathcal{B}_{k-1}\cup\{B\}$ and $\mathcal{A}_{k+1}=\mathcal{A}_{k}\cup B$.
+    Set $\mathcal{P}_{k}=\mathcal{P}_{k-1}\cup\{(F_{1},F_{2},\omega,p_{1},p_{2})\}$, $\mathcal{B}_{k}=\mathcal{B}_{k-1}\cup\{B\}$ and $\mathcal{A}_{k+1}=\mathcal{A}_{k}\cup\{B\}$.
   5. (5)
@@
 Figure 2. Step 2 of the proof of Proposition 4.8: Finding the ball in $\mathcal{B}_{k}$ that corresponds to $\text{Base}\,{\mathcal{A}_{k}}$.
-Since the $A_{i}$ are convex, this process will terminate after $K\leq m$ steps.
+Since the original cover $\mathcal{A}_{1}$ is finite, this process will terminate after $K\leq\#\text{Cluster}\,{\mathcal{A}}$ steps.
 Finally, define $\mathcal{A}$ by removing from $\mathcal{A}_{K}$ those $A_{i}$ with $A_{i}\subset B$ for some $B\in\mathcal{B}_{K}$.
@@
 Reindex $\mathcal{B}_{K}$ (and hence $\mathcal{P}_{K}$) such that $\textnormal{diam}B_{i}\geq\textnormal{diam}B_{j}$ for all $1\leq i\leq j\leq K$ and define
-:=$\mathcal{B}:=\{B_{i}\in\mathcal{B}_{K}:B_{i}\cap B_{j}=\emptyset\ \forall 1\leq j<i\}.$
+$\mathcal{B}:=\{B_{i}\in\mathcal{B}_{K}:B_{i}\cap B_{j}=\emptyset\ \forall 1\leq j<i\}.$
 Also define
-$\mathcal{P}:=\{(F_{1},F_{2},\pi,p_{1},p_{2})_{k}\in\mathcal{P}_{K}:B_{k}\in\mathcal{B}\}.$
+$\mathcal{P}:=\{(F_{1},F_{2},\omega,p_{1},p_{2})_{k}\in\mathcal{P}_{K}:B_{k}\in\mathcal{B}\}.$
 Then the partitions in $\mathcal{P}$ now satisfy P1., as required. We also have that $\mathcal{B}_{K}$ is covered by $3\mathcal{B}$ and thus $\mathcal{A}\cup 3\mathcal{B}$ forms a cover of $F$.
@@
     and hence
-    #$\#\text{Nbours}(B)\leq 4C^{2}.$
+    $\text{Nbours}(B)\leq 4C^{2}.$
     Thus we may connect the centre of $B$ to the centre of each ball in $\text{Nbours}(B)$ using a continuum of length at most $4C^{2}\cdot 3\hat{r}$.
@@
 Note that by construction of $B$ we necessarily have
-    $d(a_{n},a_{k})\geq\pi,\qquad\forall n\neq k$
+    $d(a_{n},a_{k})\geq\omega,\qquad\forall n\neq k$
     and
-    $B(a_{n},\pi/2)\subset B_{\boldsymbol{\underline{\hat{r}+\pi/2}}}.$
+    $B(a_{n},\omega/2)\subset B_{\boldsymbol{\underline{\hat{r}+\omega/2}}}.$
     Hence
-    $\frac{N}{C}\pi/2\leq\sum_{n=1}^{N}\mathcal{H}^{1}(F\cap B(a_{n},\pi/2))$ $\leq\mathcal{H}^{1}(F\cap B_{\boldsymbol{\underline{\hat{r}+\pi/2}}})$
-    $\leq C(3(2R+1)\pi/2+\hat{r}+\pi/2).$
-    Since, by contruction, $\pi\geq\hat{r}/4$, this implies
-    $N\leq C^{2}(3(2R+1)+2\hat{r}/\pi+1)\leq C^{2}(3(2R+1)+9).$
+    $\frac{N}{C}\omega/2\leq\sum_{n=1}^{N}\mathcal{H}^{1}(F\cap B(a_{n},\omega/2))$ $\leq\mathcal{H}^{1}(F\cap B_{\boldsymbol{\underline{\hat{r}+\omega/2}}})$
+    $\leq C(3(2R+1)\omega/2+\hat{r}+\omega/2).$
+    Since, by contruction, $\omega\geq\hat{r}/4$, this implies
+    $N\leq C^{2}(3(2R+1)+2\hat{r}/\omega+1)\leq C^{2}(3(2R+1)+9).$
     Connect the centre of $B$ to each $a_{n}$ to create a continuum of length at most
-    $C^{2}(3(2R+1)+9)\cdot(\pi/2+\hat{r})\leq 6C^{2}(3(2R+1)+9)\pi.$
+    $C^{2}(3(2R+1)+9)\cdot(\omega/2+\hat{r})\leq 6C^{2}(3(2R+1)+9)\omega.$
     Repeat this for every $B\in 3\mathcal{B}$ to form a union of continua $L_{3}$ which satisfies
@@
 All together we have a continuum $\tilde{L}=L\cup L_{1}\cup L_{2}\cup L_{3}$ such that
-$\mathcal{H}^{1}(\tilde{L})$ $\leq\mathcal{H}^{1}(L)+96C^{3}\mathcal{H}^{1}(F)+72C^{3}\mathcal{H}^{1}(F)+6C^{2}(R+2)(1-\varepsilon)\mathcal{H}^{1}(F)$
+[…]
 $<\mathcal{H}^{1}(L)+200C^{3}(R+2)\mathcal{H}^{1}(F).$
 Thus, (4.11) gives
-$\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
+[…]
 Since each element of $\mathcal{A}$ has diameter $\hat{r}<r$,
-$\mathcal{H}^{1}(F\cap\bigcup\mathcal{A})\leq\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
+[…]
 Therefore, since $\mathcal{A}\cup 3\mathcal{B}$ covers $F$,
@@
 S2. $\quad\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,x)\subset\mathcal{Q}_{*}(1/6,R,1/\hat{d})\cap\mathcal{C}^{*}.$
-By the virtue of the fact that $Q_{*}$ are nested, $\mathcal{Q}_{*}(\delta,R,C)\subset\mathcal{Q}_{*}(\delta,R+1,C)$, we can assume that $R$ is such that $128/(R+1)<\hat{d}/$ 12.
+By the virtue of the fact that $Q_{*}$ are nested, $\mathcal{Q}_{*}(\delta,R,C)\subset\mathcal{Q}_{*}(\delta,R+1,C)$, we can assume that $R$ is such that $8/(R+1)<\hat{d}/12$.
 We may then find a further subset $G\subset F$ with
@@
 […]
-P3.$\quad\sum_{i}\pi_{i}>\frac{1-\varepsilon}{3(2R+3)}\mathcal{H}^{1}(G).$
-For a moment fix a $P=(G_{1},G_{2},\hat{\pi},p_{1},p_{2})$ in $\mathcal{{P}}$ and set $\pi=2\hat{\pi}$.
-Because $\pi<r_{1}$, we can find $(\nu,\rho,y)\in\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,p_{1})$ such that (5.3) holds for $r=\pi$.
-Then by Theorem 2.26, there exists compact $p_{1}\in K_{F}\subset F\cap B(p_{1},\pi(R+1))$ and $y\in K_{\nu}\subset\operatorname{supp}\nu\cap B(y,R+1)$ such that
-M1. $\max\left\{\frac{\mathcal{H}^{1}(F\cap B(p_{1},\pi(R+1))\backslash K_{F})}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},\nu(B(y,R+1)\backslash K_{\nu})\right\}<\frac{1}{R+1},$
-M2. $d_{pmGH}\left((K_{F},\frac{\mathcal{H}^{1}\llcorner K_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},p_{1}),(K_{\nu},\nu\llcorner K_{\nu},y)\right)<\frac{3}{R+1}.$
+[…]
+For a moment fix a $P=(G_{1},G_{2},\hat{\omega},p_{1},p_{2})$ in $\mathcal{{P}}$ and set $\omega=2\hat{\omega}$.
+Because $\omega<r_{1}$, we can find $(\nu,\rho,y)\in\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,p_{1})$ such that (5.4) holds for $r=\omega$.
+Then by Theorem 2.29, there exists compact $p_{1}\in K_{F}\subset F\cap B(p_{1},\omega(R+1))$ and $y\in K_{\nu}\subset\operatorname{supp}\nu\cap B(y,R+1)$ such that
+M1. $\max\left\{\frac{\mathcal{H}^{1}(F\cap B(p_{1},\omega(R+1))\backslash K_{F})}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))},\nu(B(y,R+1)\backslash K_{\nu})\right\}<\frac{1}{R+1},$
+M2. $d_{pmGH}\left((K_{F},\frac{\mathcal{H}^{1}\llcorner K_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))},p_{1}),(K_{\nu},\nu\llcorner K_{\nu},y)\right)<\frac{3}{R+1}.$
 By M2. and Proposition 2.30, there exists a metric space $(Z,\lambda)$ and isometric embeddings
@@
 such that
-(5.4) $F_{z}\left(\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},\bar{\nu}\llcorner\bar{K}_{\nu}\right)+d_{H(z)}(\bar{K}_{F},\bar{K}_{\nu})<\frac{3}{R+1},$
+(5.5) $F_{z}\left(\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))},\bar{\nu}\llcorner\bar{K}_{\nu}\right)+d_{H(z)}(\bar{K}_{F},\bar{K}_{\nu})<\frac{3}{R+1},$
 where for clarity and ease of notation, the image of any set or point under these embeddings is denoted with a bar.
 We now demonstrate the existence of points in both $\bar{K}_{F}$ and $\bar{K}_{\mu}$ near $\bar{p}_{2}$.
-Firstly, note that by our choice of $R$ we now satisfy the assumptions of Proposition 3.4 where $\varepsilon=3/(R+1)$ and $\delta=1/6$, thus there exists $s\in K_{F}\cap B_{d/\pi}(p_{2},1/6)$.
+Firstly, note that by our choice of $R$ we now satisfy the assumptions of Proposition 3.5 where $\varepsilon=3/(R+1)$ and $\delta=1/6$, thus there exists $s\in K_{F}\cap B_{d/\omega}(p_{2},1/6)$.
 This point combined with (5.5), gives us the existence of $\bar{t}\in\bar{K}_{\nu}$ such that $\lambda(\bar{s},\bar{t})\leq 3/(R+1)$.
@@
 $\lambda(z,\bar{t})$ $\leq\lambda(z,\bar{p}_{2})+\lambda(\bar{p}_{2},\bar{s})+\lambda(\bar{s},\bar{t})$
-$=d_{d/\pi}(p_{1},p_{2})+d_{d/\pi}(p_{2},s)+\lambda(s,t)$
+$=d_{d/\omega}(p_{1},p_{2})+d_{d/\omega}(p_{2},s)+\lambda(s,t)$
 $\leq 1/2+1/6+3/(R+1)\leq 1.$
@@
 and $B(\bar{q},1/12)\subset B(z,R+1/24)\subset B(z,R+1)$. Thus there exists the required points near $\bar{p}_{2}$.
-Now, since $\nu\in\mathcal{A}(1/\hat{d})$,
+[…]
 (5.6) $\nu(B(q,1/12)\cap K_{\nu})\geq\frac{\hat{d}}{12}-\frac{1}{R+1}$
 which combined with (5.5) implies that
-(5.6) $\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}(B(\bar{q},1/12))}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))}\geq\nu(B(\bar{q},1/12)\cap K_{\nu})-\frac{3}{R+1}\geq\frac{\hat{d}}{12}-\frac{4}{R+1}.$
+(5.7) $\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}(B(\bar{q},1/12))}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))}\geq\nu(B(\bar{q},1/12)\cap K_{\nu})-\frac{3}{R+1}\geq\frac{\hat{d}}{12}-\frac{4}{R+1}.$
 Therefore,
-$\mathcal{H}^{1}(F\cap B_{d/\pi}(p_{1},R/2)\backslash G_{1}\cup G_{2})$ $\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\mathcal{H}^{1}(F\cap B(p_{1},\pi))$
-$\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\hat{d}\pi>\frac{\hat{d}^{2}}{24}\pi.$
-Since the Besicovitch partitions are disjoint (condition P1.), the balls $B_{d/\pi}(p_{1},R/2)$ corresponding to each partition are disjoint.
+$\mathcal{H}^{1}(F\cap B_{d/\omega}(p_{1},R/2)\backslash G_{1}\cup G_{2})$ $\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\mathcal{H}^{1}(F\cap B(p_{1},\omega))$
+$\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\hat{d}\omega>\frac{\hat{d}^{2}}{24}\omega.$
+Since the Besicovitch partitions are disjoint (condition P1.), the balls $B_{d/\omega}(p_{1},R/2)$ corresponding to each partition are disjoint.
 We may repeat the same process for each partition $P_{i}\in\mathcal{P}$ and use P3.
 to obtain
-$\mathcal{H}^{1}(F\backslash G)$ $\geq\frac{\hat{d}^{2}}{24}\cdot\frac{1-\varepsilon}{3(2R+3)}\mathcal{H}^{1}(G)$
-$=\frac{\hat{d^{2}}(1-\varepsilon)}{72(2R+3)}(\mathcal{H}^{1}(F)-\mathcal{H}^{1}(F\backslash G)).$
+[…]
+[…]
 This implies
@@
 which contradicts (5.2) and concludes the proof. ∎
-Finally, we restate the necessary part of [2, Theorem 1.2], with which we obtain our final statement.
+Finally, we restate the necessary part of [3, Theorem 1.2], with which we obtain our final statement.
 For $K\geq 1$, let $\textnormal{biLip}(K)^{*}$ be the set of all pointed metric measure spaces whose metric is $K$-bi-Lipschitz equivalent to the Euclidean norm.